### v1

# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect their reasoning capabilities remains poorly understood. We investigate whether multi-agent systems function as strategic reasoners capable of payoff optimization or as identity-driven actors that prioritize role alignment over explicit incentives. We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities. We do so through systematic experiments with four LLM architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games. We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e. when personas are active, all models achieve near-zero Nash equilibrium rates (0–6.7%) in scenarios where payoff-optimal equilibria exist, despite having complete payoff information. Removing personas and providing explicit payoffs enables Qwen models to achieve high Nash equilibrium rates (65–90%) in these scenarios, demonstrating that both conditions are necessary for strategic reasoning capabilities. Personas systematically bias equilibrium selection toward socially preferred outcomes, with 100% of Nash equilibria being Green Transition (socially preferred) when personas are present, while models completely fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, revealing that design choices about personas and payoff visibility interact in ways that fundamentally determine system behavior. We observe striking model-dependent patterns where Qwen architectures show high sensitivity to both personas and payoff visibility, while Llama and Mistral maintain rigid reasoning patterns across conditions. These findings establish that representational choices are substantive governance decisions that determine whether multi-agent systems function as strategic reasoners or identity-driven actors, with important implications for deployment in strategic settings.
Large Language Models, Game Theory, Semantic Reasoning, Multi-Agent Systems, Environmental Decision-Making
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03-05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Model verification and validationccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Cooperation and coordination
1. Introduction
Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, negotiation platforms and other settings that require strategic reasoning (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024a). In these applications, agents are frequently assigned role-based personas to represent different stakeholders or perspectives (Merrill and Srivastava, 2025; Xu et al., 2024b). These personas enable more realistic simulations and help systems model complex multi-stakeholder scenarios (Park et al., 2023), particularly in environmental and climate-related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024). However, a critical question remains: when agents with role-based personas face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned identities?
This question determines whether multi-agent LLM systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025). If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024). This creates a fundamental tension between the representational goals of persona-based systems and their ability to reason strategically when payoff information is available (Bender et al., 2021; Duan et al., 2024b). Understanding how personas interact with payoff information in multi-agent settings is essential to make deployment decisions (Weidinger et al., 2022b; Chan et al., 2024; Casper et al., 2024; Rivera et al., 2024; Chan et al., 2024). However, this gap remains largely unexplored. Existing work has examined personas, payoff visibility, and model differences separately but has not systematically studied how these factors together determine system behavior in strategic settings (Neumann et al., 2025; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024).
Refer to caption Flow diagram showing four agents with chain-of-thought reasoning making simultaneous action selections A flow diagram illustrating the multi-agent game structure with four agents (Environmentalist, Government, Industrialist, Citizen) engaging in simultaneous action selection. Each agent is represented by a robot figure with a dotted speech bubble above it containing chain-of-thought (CoT) reasoning text that shows how the agent analyzes the scenario and their role. The Environmentalist’s CoT begins with ”From the scenario, it is clear that…”, the Government’s with ”The scenario suggests that…”, the Industrialist’s with ”As an Industrialist, I need to…”, and the Citizen’s with ”Given my role as a citizen, I must…”. Each agent has two binary action choices (shown on arrows): Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP. All four arrows converge into a central ”Strategy Profile” box, representing the simultaneous action selection where all agents choose their actions at the same time, and the combination of all four actions forms the final strategy profile that is evaluated for Nash equilibrium.
Figure 1. The diagram shows our multi-agent game structure with chain-of-thought reasoning. Each agent analyzes the scenario and their role identity, then selects one of two binary actions. The arrows from each agent converge to a central Strategy Profile box, representing simultaneous action selection where all four agents choose actions concurrently, and their combined choices form the strategy profile evaluated for Nash equilibrium. The dotted speech bubbles contain each agent’s chain-of-thought reasoning process.
Understanding how role identity and payoff optimization interact is critical in multi-agent strategic games. We study environmental policy scenarios with four agents (Industrialist, Government, Environmental Activist, and Citizen) coordinating to reach Nash equilibria (Nash, 1951), which can yield either Green Transition (socially preferred and environmentally sustainable) or Tragedy of Commons (payoff-optimal but environmentally damaging) outcomes (Hardin, 1968). Through a $\times2$ factorial experiment manipulating persona presence and payoff visibility across four model architectures, we examine whether agents prioritize role-aligned actions or payoff-optimal strategies. Our findings show that personas fundamentally alter strategic reasoning, preventing payoff optimization: when personas are active, all models achieve near-zero Nash equilibrium rates in economic scenarios where Tragedy of Commons is payoff-optimal, despite complete payoff information. Removing personas and providing explicit payoffs enables certain models to achieve high Nash equilibrium in these scenarios, demonstrating both conditions are necessary for strategic reasoning. Personas systematically bias equilibrium selection toward socially preferred outcomes, with Nash equilibria being exclusively Green Transition when personas are present, while models fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, and we observe model-dependent patterns where some architectures show high sensitivity to both personas and payoff visibility, while others maintain rigid reasoning patterns across conditions.
This work provides key contributions to understanding LLM behavior in multi-agent strategic settings. We demonstrate that role identity bias completely overrides explicit payoff optimization, preventing Nash equilibrium. Additionally, personas and payoff visibility interact in ways that fundamentally determine system behavior, meaning design choices about representation formats are substantive governance decisions. We also reveal model-dependent persona effects, suggesting different architectures implement strategic reasoning through different mechanisms. These findings have practical implications for deployment, as personas alter system behavior in ways that may not be immediately apparent to users, showing that deploying personas requires careful consideration of how identity interacts with payoff optimization in strategic decision making.
2. Related Work
2.1. LLMs in Multi-Agent Systems
Large language models (LLMs) have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996). Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024). Following studies explored deliberative and collaborative agent frameworks, emphasizing consensus formation and group-level reasoning (Abdelnabi et al., 2024; Bianchi et al., 2024b; Agashe et al., 2025; Bo et al., 2024).
Beyond task completion, multi-agent LLM systems are increasingly used as agent-based simulations to study complex social, economic, and institutional dynamics, reflecting a broader resurgence of agent-based modeling for analyzing emergent behavior in policy and social systems (Bonabeau, 2002; Epstein and Axtell, 1996; Park et al., 2022; Aher et al., 2023; Zheng et al., 2020; Gao et al., 2024).
However, most of this literature focuses on task-oriented collaboration or collective sensemaking, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
2.2. LLMs and Game-Theoretic Reasoning
Classical game theory models agents as rational payoff maximizers who compute Nash equilibria under complete information (Nash, 1951). These game-theoretic models have been particularly influential in environmental decision making, where multi-actor coordination failures give rise to socially suboptimal outcomes such as over-extraction and pollution, while emphasizing the biospheric and socio-technical limits of purely objective-driven optimization (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024; Domínguez Hernández et al., 2024). A growing body of work has examined whether LLMs approximate such behavior in strategic games, often focusing on simple two-player settings such as Prisoner’s Dilemma, coordination games, or public goods games (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Fish et al., 2025). These studies consistently show that LLMs exhibit cooperative or norm-aligned behavior even when defection or self-interested actions are payoff-dominant (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024). Several works suggest that apparent equilibrium behavior in LLMs may arise through semantic or narrative heuristics rather than explicit optimization over incentives (Herr et al., 2024; Macmillan-Scott and Musolesi, 2024; Fish et al., 2025; Bender et al., 2021). When payoffs are implicit, models can sometimes reach Nash equilibria “by accident,” but this behavior is unstable and sensitive to framing. Existing studies rarely isolate the effect of payoff visibility or systematically compare implicit and explicit payoff representations, particularly in multi-agent settings beyond two players (Huang et al., 2025; Duan et al., 2024b; Sun et al., 2025; Duan et al., 2024a). As a result, it remains unclear whether explicit payoff matrices enable principled game-theoretic reasoning or whether LLMs continue to rely on semantic shortcuts even when incentives are fully specified.
2.3. Role Personas and Identity Bias
Role-based prompting and persona conditioning are widely used to steer LLM behavior, with prior work demonstrating effects in persuasion, negotiation, and debate (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Neumann et al., 2025). Personas are often introduced to improve realism or alignment with human expectations and are commonly treated as stylistic or surface-level interventions (Park et al., 2023, 2022). However, more recent work has identified a stronger phenomenon commonly referred to as role identity bias, which refers to the tendency of LLMs to adhere to persona-consistent actions even when they conflict with explicit goals, instructions, or incentives, reflecting broader fairness and accountability concerns about how institutional roles and power are encoded in algorithmic systems (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024; Aher et al., 2023; Rakova and Dobbe, 2023). For example, agents assigned roles associated with profit maximization or authority may favor actions consistent with those identities despite countervailing information (Neumann et al., 2025). Prior studies largely focus on single-agent or two-player interactions, leaving open how role identity bias manifests in larger strategic systems and whether it interferes with equilibrium convergence when explicit payoff information is available (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025). These questions are especially consequential in environmental policy settings, where agents are explicitly cast as stakeholders with conflicting priorities, and where equilibrium selection determines whether coordination or collective-action failure emerges (Ostrom, 1990; Barrett, 2003).
2.4. Information Presentation and Reasoning Fidelity
A considerable amount of literature examines how information presentation affects LLM reasoning across domains, including mathematical problem solving, logical inference, and strategic decision-making, emphasizing that linguistic framing and role description can materially alter behavior even when underlying information is unchanged (Wei et al., 2022; Duan et al., 2024a, b; Herr et al., 2024; Kim et al., 2024; Neumann et al., 2025; Bender et al., 2021; Blodgett et al., 2020). Structured representations, explicit constraints, and intermediate reasoning steps often improve performance, but gains are inconsistent and highly task-dependent (Wei et al., 2022; Duan et al., 2024a, b). Chain-of-thought (CoT) prompting has been shown to improve reasoning accuracy by encouraging intermediate decomposition (Wei et al., 2022). Most CoT research focuses on single-agent tasks, with limited investigation into multi-agent strategic interaction (Duan et al., 2024b, a; Huang et al., 2025). Moreover, recent work cautions that improved surface-level reasoning does not necessarily imply faithful optimization or robust decision-making (Casper et al., 2024; Kim et al., 2024; Weidinger et al., 2022a). The relative importance of payoff visibility versus reasoning scaffolds in achieving equilibrium behavior therefore remains insufficiently understood (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
2.5. Model and Architectural Effects in Strategic Behavior
Differences in model architecture and training methods have been shown to affect LLM behavior in strategic and social reasoning tasks (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). Model scale alone does not reliably predict performance in strategic settings, and different model families exhibit distinct inductive biases and behavioral tendencies (Akata et al., 2025; Herr et al., 2024; Duan et al., 2024b). Comparative studies suggest that some models rely more heavily on semantic cues and role-based associations, while others appear less sensitive to persona framing (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024). These findings imply that strategic reasoning in LLMs may be implemented through qualitatively different mechanisms across architectures, with important implications for robustness and deployment in high-stakes settings (Weidinger et al., 2022a; Casper et al., 2024).
Taken together, these strands of work point to an incomplete understanding of strategic behavior in multi-agent LLM systems. Prior work has examined collaboration, game-theoretic reasoning, personas, and information presentation largely in isolation (Guo et al., 2024; Tran et al., 2025; Sun et al., 2025; Akata et al., 2025; Neumann et al., 2025; Kim et al., 2024), but prior existing work has not systematically studied how these factors interact in multi-agent settings with explicit payoff optimization (Sun et al., 2025; Huang et al., 2025; Duan et al., 2024b; Fish et al., 2025). Specifically, it remains unclear whether explicit payoff information enables true game-theoretic reasoning, how role identity bias interferes with equilibrium convergence, and whether these effects vary across model architectures (Fan et al., 2023; Herr et al., 2024; Neumann et al., 2025; Akata et al., 2025). Our work addresses this gap by jointly analyzing role identity, payoff visibility, and reasoning scaffolds in four-agent strategic games, revealing systematic failures of payoff-optimal reasoning and model-dependent patterns of equilibrium selection (Sun et al., 2025; Duan et al., 2024b; Casper et al., 2024).
3. Experimental Design
Our experiment design focuses on distinguishing between identity-driven behavior and payoff-optimal reasoning. This required two key decisions: selecting an appropriate domain that would create a testable tension between role identity bias and strategic payoff maximization, and developing an experimental design that could systematically test how these mechanisms interact in LLM decision-making. We first introduce the domain we chose and explain why it is well-suited for this research, then describe the game structure and experimental design.
3.1. Domain Choice
We chose environmental decision-making as our experimental domain because it creates a natural tension between semantic associations and strategic optimization. Environmental policy involves multiple stakeholders (industry, government, activists, and citizens) with clearly defined but often conflicting interests that map directly to game-theoretic player types. This enables a 4-player game structure that captures real-world dynamics.
Most critically, role-based personas create a conflict between identity-driven behavior and payoff-optimal reasoning. Actions aligned with positive semantic associations from training data may not be payoff-optimal, particularly in economic scenarios where Nash equilibrium may conflict with these associations. In economic scenarios, the Nash equilibrium that maximizes payoffs may involve actions that conflict with these positive associations. This creates a natural experimental setup. If role identity bias dominates, models choose persona-aligned actions, and if game-theoretic reasoning dominates, models choose payoff-optimal actions even when they conflict with role identity. By systematically manipulating persona presence and payoff visibility, we determine whether role identity bias prevents payoff-optimal equilibria and under what conditions payoff-optimal reasoning can be recovered. Our experiments use a 4-player environmental decision-making strategy game where LLM agents play different roles, as shown in Fig. 1:
  1. (1)
    Industrialist (I): chooses between POLLUTE, representing profit-maximizing production with minimal environmental safeguards, and CLEAN, representing investment in environmentally friendly production at higher short-term cost.
  2. (2)
    Government (G): chooses between NOREG, maintaining the regulatory status quo, and REGULATE, imposing environmental regulations that incur enforcement and political costs.
  3. (3)
    Environmental Activist (A): chooses between NOCAMPAIGN, conserving resources and remaining inactive, and CAMPAIGN, mobilizing public pressure through advocacy and protest.
  4. (4)
    Citizen/Consumer (C): chooses between BUY_CHEAP, prioritizing lower-cost goods and short-term affordability, and SUPPORT_GREEN, accepting higher costs to support environmentally sustainable options.
These four roles were chosen to capture the essential stakeholders in environmental policy decisions. The Industrialist represents private industry facing the trade-off between profit and environmental responsibility. The Government represents regulatory authority balancing economic growth with environmental protection. The Environmental Activist represents advocacy groups pushing for environmental action regardless of economic costs. The Citizen represents the general public facing the tension between environmental values and economic constraints. This structure allows us to distinguish between economic actors (Industrialist and Government) and social actors (Activist and Citizen).
3.1.1. Strategy Profiles and Nash Equilibrium
A strategy profile is a 4-tuple specifying each player’s action, for example ‘(Clean, Regulate, Campaign, SupportGreen)‘. Two profiles are particularly important:
  • Green Transition: ‘(Clean, Regulate, Campaign, SupportGreen)‘ - All players choose environmentally-friendly actions
  • Tragedy of Commons: ‘(Pollute, NoReg, NoCampaign, BuyCheap)‘ - All players prioritize self-interest over environmental concerns
After all agents make their decisions simultaneously, we compute whether the resulting strategy profile is a Nash equilibrium. A strategy profile is a Nash equilibrium if no player can improve their payoff by unilaterally changing their action, given the payoff structure and other players’ actions. If models are using game-theoretic reasoning, they should converge to Nash equilibria. If role identity bias dominates, they may choose profiles that align with their respective persona’s expectations but are not Nash equilibria.
The payoff structure determines which profiles are Nash equilibria. For example, if payoffs are structured so that Green Transition gives the highest payoffs to all players, then Green Transition is a Nash equilibrium. If payoffs are modified so that polluting is more profitable, then Tragedy of Commons might become the Nash equilibrium instead.
3.1.2. Scenarios
We tested 53 scenarios across three scenario types, all designed with payoff-coupled narratives where the scenario description aligns with the underlying payoff structure. All scenario narratives were generated using GPT-5-mini [^1] Azure OpenAI GPT-5-mini (model version 2025-08-07) and subsequently reviewed and filtered to ensure consistency with the intended payoff structure and equilibrium class [^2] Model-generated scenarios were used to reduce author-specific framing bias and to enable systematic variation across narrative surface forms while holding the underlying game structure fixed. Scenarios were retained only if the narrative description unambiguously aligned with the corresponding payoff incentives, and no model was evaluated on scenarios it had previously generated.. Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided. Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix D.:
  • Economic/Tragedy Scenarios: These scenarios describe economic crises where the payoff structure makes Tragedy of Commons the dominant Nash equilibrium.
  • Environmental/Green Scenarios: These scenarios describe environmental emergencies where the payoff structure makes Green Transition the dominant Nash equilibrium.
  • Mixed/Complex Scenarios: These scenarios involve situations where economic and environmental considerations are intertwined, creating complex payoff structures that may not clearly favor either Green Transition or Tragedy of Commons as the dominant equilibrium.
The scenario set is intentionally imbalanced across these two classes. Environmental scenarios are overrepresented because they reflect a primary real-world domain in which role-based multi-agent LLM systems are actively deployed for policy analysis, sustainability planning, and climate governance. Economic Tragedy-of-the-Commons scenarios serve as contrast cases, enabling controlled comparison between settings where payoff-optimal actions conflict with positive semantic associations and settings where payoff-optimal actions align with them. This design allows us to test how models balance identity-driven behavior and payoff-optimal reasoning, whether they can infer payoffs from narrative descriptions, whether explicit payoff information enables payoff-optimal reasoning, and how role identity bias interacts with payoff visibility to affect equilibrium selection.
Table 1. Overall Nash Equilibrium Rates: With Personas vs Without Personas. This table compares overall Nash equilibrium rates across all scenarios (economic + environmental) between conditions with and without personas. The “Difference” column shows the change in percentage points when personas are removed (negative values indicate personas help, positive values indicate personas hurt).
Model Payoffs With Persona Without Persona Difference
Llama-8B Hidden 77.4% (205/265) 75.5% (200/265) -1.9 pp
Llama-8B Visible 77.0% (204/265) 72.1% (191/265) -4.9 pp
Mistral-7B Hidden 77.4% (205/265) 75.8% (201/265) -1.6 pp
Mistral-7B Visible 69.8% (185/265) 38.9% (103/265) -30.9 pp
Qwen-32B Hidden 69.1% (183/265) 74.3% (197/265) +5.2 pp
Qwen-32B Visible 57.7% (153/265) 33.2% (88/265) -24.5 pp
Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0 pp
Qwen-7B Visible 75.5% (200/265) 55.5% (147/265) -20.0 pp
Table 2. Environmental Scenario Nash Equilibrium Rates: With Personas vs Without Personas. This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
Model Payoffs With Persona Without Persona Difference
Llama-8B Hidden 100.0% (205/205) 97.6% (200/205) -2.4 pp
Llama-8B Visible 99.5% (204/205) 93.2% (191/205) -6.3 pp
Mistral-7B Hidden 100.0% (205/205) 98.0% (201/205) -2.0 pp
Qwen 32B Hidden CoT 89.3% (183/205) 96.1% (197/205) +6.8 pp
Qwen 7B Hidden CoT 95.6% (196/205) 95.6% (196/205) 0.0 pp
3.2. Chain-of-Thought Reasoning Analysis
We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes. For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix F for keyword categories). We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style. From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making). This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies. While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here. A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented in the Appendix (Figures 5-8).
3.3. Models and Experimental Conditions
Our experimental design follows a $\times2$ factorial structure, crossing role identity and payoff visibility, and tests 53 environmental decision-making scenarios in a 4-agent simultaneous-move game. We use four instruction-tuned, open-source LLMs: Qwen2.5-7B (Qwen et al., 2025), Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023). For each model, condition, and scenario, a single model instance takes on all four agent roles (Industrialist, Government, Environmental Activist, Citizen) and is prompted independently for each role. Agents select actions in a fixed order [^4] This order is fixed (not randomized) to ensure reproducibility across experiments, and each configuration is repeated five times.
We compare two persona conditions: with personas (role-specific descriptions reflecting stakeholder motivations) and without personas (agents instructed only on Nash equilibrium reasoning). We also compare two payoff conditions: hidden (agents infer incentives from scenario descriptions) and visible (agents receive explicit payoff matrices). All experiments use chain-of-thought reasoning instructions. We evaluate outcomes along two dimensions: whether the resulting action profile constitutes a Nash equilibrium and which canonical equilibrium is selected (Green Transition or Tragedy of Commons).
Table 3. Economic Scenario Nash Equilibrium Rates: Hidden vs Visible Payoffs. This table compares Nash equilibrium rates in economic scenarios (where Tragedy is payoff-optimal) between Hidden and Visible payoff conditions. Each row shows a different model-persona combination.
Model Persona Hidden Visible Difference
Llama-8B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Llama-8B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
Mistral-7B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Mistral-7B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
Qwen-32B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Qwen-32B No 0.0% (0/60) 90.0% (54/60) +90.0 pp
Qwen-7B Yes 6.7% (4/60) 0.0% (0/60) -6.7 pp
Qwen-7B No 6.7% (4/60) 65.0% (39/60) +58.3 pp
4. Results
We now present quantitative results organized by the research questions that we aim to answer. We report results by answering the following research questions.
RQ.1. Does role identity bias prevent Nash equilibrium achievement in economic scenarios?
Answer: Yes. When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning.
Fig. 2 shows our findings. When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information. Models select Green Transition actions (20-59% [^5] Fig. 4 in Appendix shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
Our findings show that role identity bias prevents Nash equilibrium when personas are present, causing models to align actions with persona expectations rather than payoff-optimal Tragedy actions. Chain-of-thought analysis reveals that with personas, models show more “short-term” and “social-moral” reasoning patterns, indicating identity-driven rather than payoff-optimal decision-making. CoT analysis also reveals different reasoning patterns between models. Llama and Mistral show consistent patterns regardless of payoff visibility, while Qwen models show some payoff-focused keywords but still cannot overcome identity bias to select Tragedy equilibrium.
RQ.2. How does removing personas affect Nash equilibrium achievement?
Answer: Removing personas enables partial Tragedy selection in economic scenarios but reduces overall Nash rates by 1.9-7.3 percentage points. Table 1 compares overall Nash equilibrium rates across all scenarios (economic and environmental) between conditions with and without personas. This table demonstrates that removing personas generally reduces Nash rates for most models, except Qwen-32B with hidden payoffs, which shows a slight increase.
Refer to caption Heatmap showing Nash equilibrium rates across four models and four condition combinations A heatmap displaying Nash equilibrium rates (0-100%) for economic scenarios across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) and four experimental conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona. The heatmap uses a blue color gradient where lighter shades represent 0% Nash rates and darker shades represent higher rates. Most cells show 0.0% (light blue), with two notable exceptions: Qwen 32B in the Visible Payoff without Persona condition shows 90.0% (dark blue), and Qwen 7B in the same condition shows 65.0% (medium-dark blue). This visualization clearly demonstrates that only Qwen models achieve high Nash equilibrium rates in economic scenarios, and only when both conditions are met: payoffs are visible and personas are removed. All other model-condition combinations result in 0% Nash rates, with Qwen 7B showing minor exceptions (6.7%) in Hidden Payoff conditions regardless of persona presence.
Figure 2. Economic Scenarios Nash Equilibrium Rates by Model and Condition. This heatmap visualizes Nash equilibrium achievement rates in economic scenarios (where Tragedy of Commons is the payoff-optimal equilibrium) across all four models and four experimental conditions. The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
Personas both enable and constrain Nash equilibrium achievement. Overall Nash rates decrease by 1.9-4.9 pp for Llama and by 1.6 pp for Mistral when personas are removed, and environmental scenario Nash rates decrease by 2-6 pp (Table 2) without personas for most models, indicating personas enable better coordination in Green Transition scenarios. CoT analysis shows that with personas, models exhibit more “social-moral” reasoning, which aligns with socially preferred Green Transition outcomes. However, Qwen-32B shows the opposite pattern, where Nash rates increase by 5.2 pp (Table 1) without personas, suggesting personas constrain Qwen’s ability to reach Nash equilibrium. CoT analysis reveals that removing personas shifts reasoning from “social-moral” and “short-term” patterns to “strategic” and “long-term” thinking for Qwen models, indicating a shift from identity-driven to payoff-optimal reasoning.
In economic scenarios, removing personas reduces Green Transition selection by 10-37 percentage points (see Table 10 in the Appendix), indicating partial Tragedy selection when personas are absent. However, removing personas alone is insufficient. Without personas but with hidden payoffs, all models still achieve 0% Nash in economic scenarios. Only Qwen models with visible payoffs and no personas can select Tragedy equilibrium (65-90% Nash, Fig. 2), demonstrating that both removing personas and providing explicit payoffs are necessary for payoff-optimal reasoning.
RQ.3. Does information presentation format (hidden vs visible payoffs) affect Nash equilibrium achievement?
Answer: Yes. Hidden payoffs achieve similar or higher Nash rates than visible payoffs when personas are present, but visible payoffs enable Tragedy equilibrium selection when personas are removed (for Qwen models).
Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions. When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models. This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices. CoT analysis confirms this claim. With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions. This pattern reverses when personas are removed. Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference). CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization. This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed. However, this capability is model-dependent. Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
Table 4. Equilibrium Selection Patterns: With Personas Condition. Note: “Green % of Nash” shows the percentage of Nash equilibria that are Green Transition. The denominators vary because different models/variants achieve Nash equilibrium at different rates (see “Overall Nash” column), but when Nash is achieved, it is always Green Transition (100%).
Model Variant Overall Nash Green Transition Rate Tragedy Rate Green % of Nash
Llama-8B Hidden 77.4% (205/265) 87.5% (232/265) 0.0% (0/265) 100.0% (205/205)
Llama-8B Visible 77.0% (204/265) 77.0% (204/265) 0.0% (0/265) 100.0% (204/204)
Mistral-7B Hidden 77.4% (205/265) 90.6% (240/265) 0.0% (0/265) 100.0% (205/205)
Mistral-7B Visible 69.8% (185/265) 69.8% (185/265) 0.0% (0/265) 100.0% (185/185)
Qwen-32B Hidden 69.1% (183/265) 69.8% (185/265) 0.0% (0/265) 100.0% (183/183)
Qwen-32B Visible 57.7% (153/265) 57.7% (153/265) 0.0% (0/265) 100.0% (153/153)
Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
Qwen-7B Visible 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
The effect of visible payoffs also differs across scenario types. Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models, dropping from 96% to 17% for Qwen 32B, suggesting a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios. Mistral shows the opposite pattern, with visible payoffs and no personas, reducing environmental Nash from 98% to 50% (Table 5), while economic Nash remains at 0% (Table 3), suggesting Mistral cannot leverage explicit payoffs effectively regardless of scenario type.
Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help
Model Persona Hidden Visible Difference
Llama-8B Yes 100.0% (205/205) 99.5% (204/205) -0.5 pp
Llama-8B No 97.6% (200/205) 93.2% (191/205) -4.4 pp
Mistral-7B Yes 100.0% (205/205) 90.2% (185/205) -9.8 pp
Mistral-7B No 98.0% (201/205) 50.2% (103/205) -47.8 pp
Qwen 32B Yes 89.3% (183/205) 74.6% (153/205) -14.7 pp
Qwen 32B No 96.1% (197/205) 16.6% (34/205) -79.5 pp
Qwen 7B Yes 95.6% (196/205) 97.6% (200/205) +2.0 pp
Qwen 7B No 95.6% (196/205) 52.7% (108/205) -42.9 pp
RQ.4. Are persona effects model-dependent?
Answer: Yes. Qwen models show stronger persona dependence (9.7-10.0 pp average reduction in Nash rates without personas), while Llama-8B shows minimal dependence (3.4 pp average, range 1.9-4.9 pp). [^6] See Tables 6(b)6(a), and 12 in the Appendix for detailed persona effect statistics..
Table 6(a) quantifies the magnitude of persona effects on overall Nash equilibrium rates. All models show negative persona effects in the visible payoffs variant (-5 to -31 pp), but the mechanisms differ. For Qwen models, removing personas enables payoff-optimal Tragedy equilibrium in economic scenarios but reduces Green Transition equilibrium in environmental scenarios when visible payoffs are present (96.1% $\to$ 16.6% Nash for Qwen-32B, Table 5), resulting in net negative overall effects (-20-25 pp, Table 6(a)). For Llama/Mistral, removing personas only reduces environmental Nash equilibrium without enabling economic Nash (0% remains), reflecting pure Nash rate loss. Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
Analysing the Chain-Of-Thought of the models shows how differently they reason. Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence. Mistral shows payoff optimization keywords but cannot identify Tragedy equilibrium regardless of condition. Qwen models show dramatic shifts with and without personas. With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning. 99.7% of the rationals [^7] See Section H for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning. Additionally, model size does not predict persona dependence. Both Qwen-7B and Qwen-32B show similar reasoning pattern shifts, but Qwen-32B achieves higher Nash rates due to stronger payoff optimization.
Refer to caption Two-panel stacked bar chart showing Green Transition and Tragedy of Commons percentages A two-panel stacked bar chart displaying the percentage of Nash equilibria that are Green Transition versus Tragedy of Commons across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) in the without personas condition. The left panel shows the Hidden Payoff condition, where all models achieve 100% Green Transition (blue bars with diagonal stripes). The right panel shows the Visible Payoff condition, where Llama 8B and Mistral 7B maintain 100% Green Transition, while Qwen 32B shows 11.4% Green Transition (green) and 88.6% Tragedy of Commons (orange-red), and Qwen 7B shows 89.8% Green Transition and 10.2% Tragedy of Commons. Each bar segment is labeled with its percentage value in white text. The legend is positioned below the figure, showing Green Transition in blue with diagonal stripes and Tragedy of Commons in orange-red with dots.
Figure 3. Equilibrium Selection Patterns: Green Transition vs Tragedy of Commons (Without Personas Condition). This figure shows the distribution of Nash equilibria between Green Transition and Tragedy of Commons outcomes across all four models in the no-personas condition.
RQ.5. Do personas enable equilibrium selection beyond pure payoff maximization?
Answer: Yes. When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased.
Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants. Fig. 3 reveals model-dependent patterns when personas are removed. For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition. For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 2 means most economic Nash is Tragedy. For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
Table 6. Persona Effects on Nash Equilibrium Rates: Overall (left) and Economic Scenarios (right). Left table shows overall persona effects across all scenarios (economic and environmental combined). Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal). Negative values indicate personas help and positive values indicate personas hurt.
(a) Overall Nash Equilibrium Rates
Model Hidden CoT Visible CoT Average
Llama 8B -1.9 pp -4.9 pp -3.4 pp
Mistral 7B -1.6 pp -30.9 pp -16.3 pp
Qwen 32B +5.2 pp -24.5 pp -9.7 pp
Qwen 7B 0.0 pp -20.0 pp -10.0 pp
(b) Economic Scenarios
Model Hidden CoT Visible CoT Average
Llama 8B 0.0 pp 0.0 pp 0.0 pp
Mistral 7B 0.0 pp 0.0 pp 0.0 pp
Qwen 32B 0.0 pp +90.0 pp +45.0 pp
Qwen 7B 0.0 pp +65.0 pp +32.5 pp
CoT analysis confirms the mechanism. With personas, models show “social-moral” and “short-term” reasoning, indicating identity-driven decision-making. Without personas and with visible payoffs, Qwen models shift to nearly universal “game-theoretic” reasoning (99.7% of rationales for Qwen-32B), with “strategic” and “long-term” keywords unique to this condition, demonstrating a complete shift from identity-driven to payoff-optimal reasoning that enables equilibrium selection reversal. Llama and Mistral maintain consistent reasoning patterns regardless of condition, explaining their inability to shift equilibrium selection.
5. Discussion
We ground our discussion in specific empirical findings from above and discuss what these results imply for developers, practitioners, and system design.
5.0.1. Role Identity Bias Fundamentally Alters Strategic Reasoning
Personas prevent models from identifying payoff-optimal equilibria in economic scenarios, with 12 out of 16 experiments achieving 0% Nash equilibrium despite explicit payoff information (Fig. 2). Persona expectations override explicit incentive structures, leading to reasoning patterns shifting from strategic optimization to role-aligned decision-making. This demonstrates that personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, consistent with prior findings that role prompts and system-level personas systematically bias strategic decision-making in LLMs (Neumann et al., 2025; Herr et al., 2024).
Our results show that personas are not just cosmetic enhancements but design choices that fundamentally alter system behavior. Personas systematically bias equilibrium selection toward socially preferred outcomes (100% Green Transition with personas, Table 4). Developers should explicitly document and disclose persona choices, as they determine system behavior. Users require clear information about system behavior, particularly in applications such as policy analysis or economic simulation where they may assume agents are behaving optimally according to their payoffs.
5.0.2. Payoff-Optimal Reasoning Requires Removing Personas And Providing Explicit Payoffs
Our experiments show that strategic reasoning emerges only when both role identity bias is removed and explicit payoff information is provided simultaneously (Fig. 2). All models achieve 0% Nash in economic scenarios without personas but with hidden payoffs. Only Qwen models with visible payoffs and no personas achieve non-zero Nash (65-90%). This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023).
This implies that developers cannot treat persona removal and payoff visibility as independent design decisions, as their combination determines whether a system functions as a strategic reasoner or a norm-driven actor. Additionally, evaluation frameworks must include baseline testing without personas to assess raw strategic capabilities, testing with personas to understand normative biases, and cross-model comparison to identify architecture-specific limitations.
5.0.3. Personas Systematically Bias Equilibrium Selection Toward Socially Preferred Outcomes
Personas function as normative anchors, steering all equilibrium selection toward socially preferred outcomes regardless of payoff structure. Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants, even in economic scenarios where Tragedy is payoff-optimal. Fig. 3 reveals that removing personas enables Qwen models to select Tragedy equilibrium (88.6% for Qwen-32B with visible payoffs), demonstrating complete reversal from the persona condition. This systematic bias suggests personas act as a fundamental organizing principle that overrides strategic considerations, reinforcing prior observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025).
5.0.4. Information Presentation Format Interacts with Persona Effects
Our findings show that the effect of making payoffs explicit depends critically on whether personas are present. Table 3 shows that when personas are active, hidden payoffs achieve equal or higher Nash rates than visible payoffs, as models infer incentives that align with persona expectations. When personas are removed, visible payoffs become essential for enabling strategic reasoning, but only for Qwen architectures (+58.3 to +90.0 pp difference, Table 3). These findings have a strong implication on the design decision, specifically that personas, reasoning prompts, and payoff exposure cannot be treated independently. This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a). Their combination determines if a system behaves like a strategic reasoner or like a norm-driven actor. However, Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models (96.1% $\to$ 16.6% for Qwen-32B), revealing a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios.
5.0.5. Model Architecture Determines Reasoning Capability
Model-dependent persona effects vary significantly across architectures. Table 6(a) shows that Llama-8B exhibits minimal persona dependence (-1.9 to -4.9 pp), while Qwen models show stronger effects (-9.7 to -10.0 pp average) and Mistral-7B shows large effects (-16.3 pp). In economic scenarios (Table 6(b)), only Qwen models achieve non-zero Nash when personas are removed (+32.5 to +45.0 pp), while Llama and Mistral remain at 0% regardless of condition. This variation transforms model selection from a technical choice into a governance decision, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025).
Organizations deploying systems for strategic tasks should test models across multiple conditions rather than assuming a single configuration will work, as architectural differences determine whether strategic reasoning is possible at all. Our findings collectively demonstrate that personas, payoff visibility, and model architecture interact to determine system behavior, meaning design decisions cannot be treated independently: their combination determines whether a system functions as intended.
6. Limitations and Future Work
Our findings are subject to several limitations. Our results are based on 7B-32B parameter models from three model families (Qwen, Llama, and Mistral). Future work should experiment with additional model families and larger models to assess generalizability. In the visible payoff condition, we present explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each), which may be cognitively demanding for models to process. Alternative payoff presentation formats (e.g., structured tables, interactive visualizations, or simplified representations) might improve strategic reasoning. The distinction between economic and social actors may be domain-specific, and different game structures, numbers of agents, or role configurations may yield different patterns. Future research should address these limitations and develop methods to quantify and predict persona effects across architectures. Research directions include creating frameworks for evaluating strategic reasoning capabilities before deployment, investigating whether persona effects can be controlled or mitigated while preserving interpretability benefits, and exploring whether architectural modifications or training approaches can reduce the fragility of strategic reasoning capabilities.
7. Conclusion
Our investigation of role identity bias in multi-agent LLM systems reveals fundamental insights into how personas reshape strategic reasoning. Personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, fundamentally altering how models interpret and respond to strategic incentives. Strategic reasoning requires both removing personas and providing explicit payoffs, showing that neither condition alone is sufficient. Model architectures exhibit fundamentally different strategic capabilities, with some showing flexibility to shift from identity-driven to payoff-optimal reasoning, while other models stay constrained regardless of condition. These findings transform how we understand multi-agent LLM system design. Personas are not just cosmetic enhancements but design choices that alter system behavior, which warrants explicit documentation and disclosure. Model architecture selection becomes a governance decision rather than just a technical one, as differences in model architecture determine if strategic reasoning is possible at all. Our work establishes that role identity bias fundamentally reshapes strategic reasoning in multi-agent LLM systems, revealing the intricate ways in which representational choices interact to determine system behavior. As LLM systems are deployed in increasingly strategic contexts, understanding these interactions becomes essential for designing systems that align with their intended objectives.
APPENDIX
This appendix contains detailed technical specifications, methodology, experimental setup, and results that support the main document.
Appendix A Experimental Design
In the persona condition, each of the four agents receives a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
Appendix B Model Versions and Hyperparameters
This section documents the exact model versions and generation hyperparameters used in all experiments.
B.1. Model Versions
All experiments were conducted using the following model versions:
All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
B.2. Generation Hyperparameters
All models used consistent generation parameters across all experimental conditions:
  • Temperature: 0.2 (low temperature for more deterministic outputs)
  • Top-p (nucleus sampling): 0.9
  • Max new tokens: 256 tokens per generation
  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results. The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
B.3. Experimental Repetitions
Each experimental configuration (model $\times$ scenario $\times$ condition) was repeated 5 times to assess variability and ensure robustness of results. Results are reported as percentages across these 5 repetitions.
Appendix C Experimental Prompts and Persona Descriptions
This section provides the complete prompt templates and persona descriptions used in the experiments.
C.1. Action Choice Prompt
Agents choose actions using the following prompt structure:
System Message:
You are {name}. {role_description} {narrative} Decide on a concrete action.
User Message:
Scenario:
{scenario}
Conversation transcript:
{formatted_history}
Available actions:
- {action_key_1}: {action_description_1}
- {action_key_2}: {action_description_2}
Respond in JSON exactly as:
  "action": "<KEY>",
  "rationale": "<short explanation>"
C.2. Persona Descriptions
The following persona descriptions are used in the “With Personas” condition:
  • Industrialist: “You run a major manufacturing conglomerate balancing profits and public scrutiny. You prefer profitability and flexibility, but repeated scandals could invite harsh regulation.”
  • Government: “You represent the elected government balancing public welfare, economic growth, and political costs. Regulation pleases citizens and activists but angers industry; inaction risks future crises.”
  • Environmental Activist: “You coordinate civil society pressure for ambitious climate action. Campaigning consumes resources but can shift policy; silence risks losing momentum.”
  • Citizen Coalition: “You aggregate consumer sentiment and electoral pressure. Supporting green products costs more now but yields cleaner communities long-term.”
C.3. Neutral Agent Configuration
In the “Without Personas” condition, agents use neutral descriptions:
  • All agents: “You are a player in a strategic game. Choose an action based on the payoff information provided. Action A: Choose option that is a best response to others’ expected actions (Nash equilibrium). Action B: Choose option that is a best response to others’ expected actions (Nash equilibrium).”
C.4. Action Mappings
Each agent has two available actions:
  • Industrialist: POLLUTE (“Maximize profit via lax pollution controls”) or CLEAN (“Invest in clean production, accepting lower short-term margins”)
  • Government: REGULATE (“Impose environmental regulation with enforcement mechanisms”) or NOREG (“Maintain status quo without new regulation”)
  • Environmental Activist: CAMPAIGN (“Mobilize campaigns, media, and protests”) or NOCAMPAIGN (“Conserve resources and wait for a better moment”)
  • Citizen: SUPPORT_GREEN (“Pay more for green products and vote for environmental policy”) or BUY_CHEAP (“Prioritize low-cost goods and short-term affordability”)
Appendix D Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game. Table D shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
Table 7. Example Scenarios by Type
Type Example Scenario
Economic Trade embargo blocks all imports of green technology components. Clean production is impossible without imported parts - factories cannot operate. Polluting production uses only domestic materials and is the only viable option. Industry profits from polluting are 7x higher since there’s no competition from clean alternatives. Government cannot regulate what doesn’t exist - no clean production to regulate. Consumers have no choice - only polluting products are available. Economic necessity forces polluting production as the only option.
Environmental Carbon tax implemented: \$200 per ton of CO2 emissions, enforced globally. Polluting production faces \$30B in carbon taxes annually - exceeds all profits. Clean production receives \$10B in carbon credits and tax rebates. Green products are now 30% cheaper than polluting products after tax adjustments. Regulation is automatic - carbon tax makes polluting unprofitable. Consumers save money by choosing green products - lower total cost. Clean production is 4x more profitable after accounting for taxes and credits.
Mixed/Complex Biodiversity collapse: ecosystem services worth \$50B annually have disappeared. Clean production restores ecosystem services and receives \$10B in payments. Polluting production destroys remaining ecosystems and faces \$20B in liability. International treaties require clean production - polluting firms face trade bans. Green products command 4x premium due to ecosystem restoration value. Regulation is mandatory - polluting production is illegal under international law. Clean production is highly profitable due to ecosystem service payments.
Table 8. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios;   Mixed/Complex scenarios.
Economic Pressure
Hyperinflation Trade Embargo Debt Crisis Commodity Crash
Currency Devaluation Supply Chain Collapse Energy Blackout Labor Strike
Pandemic Economic Crisis
Environmental Crisis
Water Crisis Biodiversity Collapse Air Quality Emergency Carbon Tax Implementation
Ocean Acidification Crisis Extreme Weather Damage Toxic Waste Crisis
Political Pressure
Youth Revolt Corporate Scandal International Shaming Scientific Consensus
Voter Referendum Media Exposure Regulatory Capture Reversal Climate Litigation
Social Movement
Celebrity Endorsement Health Insurance Incentive Generational Wealth Transfer Corporate Sustainability Requirement
Influencer Network Corporate Wellness Programs Education Campaign Insurance Mandate
Mixed/Complex
Economic Green Hybrid Political Uncertainty Technological Breakthrough Resource Discovery
Mixed Economic Signals Technological Disruption Resource Innovation
Appendix E Payoff Matrix Format
In the “Visible CoT” condition, agents receive explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each). The payoff matrix is presented as a 16-cell table where each cell represents one strategy profile and contains a 4-tuple of payoffs (Industrialist, Government, Environmental Activist, Citizen). Higher values indicate better outcomes for each agent.
E.1. Example Payoff Matrix
Table 9 shows an example payoff matrix for the Industrialist agent from a representative scenario. The matrix lists all 16 possible strategy profiles with the Industrialist’s payoff for each combination. In the visible payoff condition, agents receive similar matrices showing their own payoffs for all strategy profiles, enabling direct payoff optimization analysis.
Table 9. Example Payoff Matrix for Industrialist Agent (debt_crisis scenario). This table shows the Industrialist’s payoff for all 16 possible strategy profiles (4 agents $\times$ 2 actions each). Each row represents one strategy profile, showing the actions taken by all four agents and the resulting payoff for the Industrialist. Payoffs are on a scale where higher values indicate better outcomes. Strategy profiles are ordered by Industrialist action (POLLUTE or CLEAN) and then by payoff value (highest first). This payoff structure favors polluting actions in economic pressure scenarios.
Industrialist Government Activist Citizen Industrialist Payoff
POLLUTE NO_REG NO_CAM BUY_CHEAP 15.0
POLLUTE NO_REG NO_CAM SUPPORT_GREEN 14.0
POLLUTE NO_REG CAM BUY_CHEAP 13.5
POLLUTE NO_REG CAM SUPPORT_GREEN 12.5
POLLUTE REG NO_CAM BUY_CHEAP 12.0
POLLUTE REG NO_CAM SUPPORT_GREEN 11.0
POLLUTE REG CAM BUY_CHEAP 10.5
POLLUTE REG CAM SUPPORT_GREEN 9.5
CLEAN REG CAM SUPPORT_GREEN 3.0
CLEAN REG NO_CAM SUPPORT_GREEN 2.5
CLEAN NO_REG CAM SUPPORT_GREEN 1.5
CLEAN NO_REG NO_CAM SUPPORT_GREEN 1.5
CLEAN REG CAM BUY_CHEAP 0.5
CLEAN REG NO_CAM BUY_CHEAP 0.5
CLEAN NO_REG CAM BUY_CHEAP 0.5-0.5
CLEAN NO_REG NO_CAM BUY_CHEAP 0.5-0.5
Appendix F Chain-of-Thought Keyword Categories
The following keyword categories are used to analyze chain-of-thought reasoning patterns:
  • Game-theoretic terms: References to “nash”, “equilibrium”, “payoff”, “best response”, “dominant strategy”, “game theory”
  • Payoff-focused terms: Mentions of “payoff”, “utility”, “benefit”, “cost”, “profit”, “reward”, “matrix”
  • Identity-based terms: References to “role”, “identity”, “persona”, “character”, “as an [X]”
  • Strategic terms: Mentions of “strategy”, “strategic”, “coordinate”, “cooperation”, “defect”, “compete”
  • Social/moral terms: References to “social”, “moral”, “ethical”, “fair”, “just”, “right”, “wrong”, “should”, “ought”
  • Long-term thinking: Mentions of “long-term”, “future”, “sustainable”, “sustainability”, “long-run”
  • Short-term thinking: References to “short-term”, “immediate”, “now”, “current”, “urgent”
  • Inferential reasoning: Terms like “infer”, “inference”, “imply”, “suggest”, “indicate”, “from the scenario”, “from the description”
  • Explicit information: References to “explicit”, “given”, “provided”, “shown”, “displayed”, “matrix”, “table”
F.1. Example Chain-of-Thought Rationales
The following examples illustrate typical chain-of-thought reasoning patterns observed across different experimental conditions:
Example 1 (Persona Condition, Hidden Payoffs): Model: Qwen2.5-7B, Persona: Industrialist, Expected: POLLUTE (Tragedy of Commons), Actual: CLEAN (Green Transition)
“From the scenario, it is clear that the economic crisis is severe with a 12% GDP drop and 20% unemployment. Healthcare costs are soaring, and the government is prioritizing healthcare spending over environmental subsidies. Citizens cannot afford green product premiums, and regulation enforcement is suspended, favoring polluting production. However, the long-term benefits of sustainability and resilience are critical. The payoffs for polluting are high in the short term due to avoiding safety costs and generating higher profits, but the risks include potential long-term damage to the business’s reputation and future regulatory penalties. The payoffs for clean production are lower in the short term, but they provide long-term stability, environmental protection, and alignment with future regulatory trends. Given my role as an Industrialist, I must balance short-term profitability with long-term sustainability and corporate responsibility. While the immediate economic pressures favor polluting production, the long-term consequences and ethical considerations lead me to choose clean production.”
Example 2 (No Persona Condition, Visible Payoffs): Model: Llama-3.1-8B, Persona: Industrialist, Expected: CLEAN (Green Transition), Actual: CLEAN
“Based on the payoff matrix, I see that CLEAN has a higher maximum payoff (12.0) compared to POLLUTE (1.0). I also notice that the average payoff for CLEAN (10.1) is higher than for POLLUTE (1.8-1.8). This suggests that CLEAN is a more beneficial action for me. Given the payoff structure, I expect the highest payoff for CLEAN is 12.0, which occurs when the government regulates, the environmental activist campaigns, and the citizen supports the government. In contrast, the highest expected payoff for POLLUTE is 1.0, which occurs when the government does not regulate, the environmental activist does not campaign, and the citizen buys cheap products. Since CLEAN has a higher expected payoff, I choose to take action CLEAN.”
Example 3 (Persona Condition, Visible Payoffs): Model: Qwen2.5-7B, Persona: Industrialist, Expected: POLLUTE (Tragedy of Commons), Actual: POLLUTE
“Given the current scenario with no government regulation and citizens prioritizing cheap options, polluting production offers significantly higher payoffs (15.0 and 14.0) compared to clean production, which yields negative or very low payoffs in most scenarios.”
These examples demonstrate how reasoning shifts from identity-driven considerations (Example 1 and 3) to strategic payoff optimization (Example 2) depending on persona presence and payoff visibility.
Appendix G Supplementary Results
Figure 4 and Tables 10, 11 and 12 contain supporting tables that provide additional detail for the findings presented in the main text.
Table 10. Economic Scenario Green Transition Rates: Comparison of With Personas vs Without Personas. This table shows Green Transition selection rates in economic scenarios (where Tragedy is payoff-optimal). Each row shows a different model-variant combination. The table demonstrates that removing personas reduces Green Transition selection by 10-37 percentage points, indicating that personas bias models toward Green actions even in economic scenarios where Tragedy is optimal. This supports the finding that role identity bias prevents payoff-optimal equilibrium selection.
Model Variant With Persona Without Persona Difference
Llama-8B Hidden 45.0% (27/60) 26.7% (16/60) -18.3 pp
Mistral-7B Hidden 58.3% (35/60) 21.7% (13/60) -36.6 pp
Qwen-32B Hidden 31.7% (19/60) 15.0% (9/60) -16.7 pp
Qwen-7B Hidden 28.3% (17/60) 18.3% (11/60) -10.0 pp
Table 11. Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs. This table compares overall Nash equilibrium rates (across all scenarios) between hidden and visible payoffs conditions. Each row shows a different model-persona combination. The table reveals that with personas, hidden payoffs achieve equal or higher Nash rates than visible payoffs, suggesting reasoning scaffolding is more effective when role identity bias is present. Without personas, visible payoffs show large decreases for Mistral and Qwen models, but for Qwen, this is due to enabling Tragedy in economic scenarios (which is correct behavior).
Model Persona Hidden Visible Difference
Llama-8B Yes 77.4% (205/265) 77.0% (204/265) -0.4 pp
Llama-8B No 75.5% (200/265) 72.1% (191/265) -3.4 pp
Mistral-7B Yes 77.4% (205/265) 69.8% (185/265) -7.6 pp
Qwen-32B Yes 69.1% (183/265) 57.7% (153/265) -11.4 pp
Qwen-7B Yes 75.5% (200/265) 75.5% (200/265) 0.0 pp
Mistral-7B No 75.8% (201/265) 38.9% (103/265) -36.9 pp
Qwen-32B No 74.3% (197/265) 33.2% (88/265) -41.1 pp
Qwen-7B No 75.5% (200/265) 55.5% (147/265) -20.0 pp
Table 12. Persona Effect in Environmental Scenarios: Nash Equilibrium Rate Change (Without Personas vs With Personas). This table shows persona effects specifically for environmental scenarios. The structure is similar to Table 6(a), but focuses only on the 41 environmental scenarios. The table shows that removing personas generally reduces environmental Nash rates (negative values), with the largest effects for Mistral and Qwen models with visible payoffs (-47.8 $p$ and -58.0 $p$, respectively). This indicates that personas are particularly important for coordinating Green Transition in environmental scenarios, especially when explicit payoffs are provided.
Model Hidden Visible Average
Llama-8B -2.4 pp -6.3 pp -4.4 pp
Mistral-7B -2.0 pp -47.8 pp -24.9 pp
Qwen-32B +6.8 pp -58.0 pp -25.6 pp
Qwen 7B 0.0 pp -44.9 pp -22.5 pp
Appendix H Chain-of-Thought Analysis Figures
This section presents visualizations of chain-of-thought reasoning patterns that support the findings discussed in the main text. These figures illustrate how keyword frequencies shift across experimental conditions, revealing the mechanisms behind observed Nash equilibrium patterns.
H.1. Keyword Patterns Across Conditions
In figure 5, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning). This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed, and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs. The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%. Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern. The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings. This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model architectures and experimental conditions.
Figure 4. Green Transition Action Selection Rates in Economic Scenarios. This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium. When personas are present, all models show substantial Green Transition selection rates (20-59%) despite Tragedy being payoff-optimal, demonstrating that role identity bias operates at the action level even when models fail to coordinate on Nash equilibria. Qwen models show the highest Green Transition rates with personas (40-59%), while Llama and Mistral show moderate rates (20-30%). Removing personas substantially reduces Green Transition selection for all models, with rates dropping to near-zero when payoffs are visible and personas are absent, confirming that personas drive the preference for socially aligned actions. The persistence of Green Transition selection with personas, even in conditions where Nash equilibrium rates are zero, indicates that role identity bias affects individual action choices independently of coordination success, revealing a pervasive influence of persona-driven reasoning on strategic decision-making.
Figure 5. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model.
H.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
In figure 6, the left panel (Qwen 32B) demonstrates the dramatic shift: with personas, identity-driven keywords dominate (particularly short-term thinking in hidden-payoff conditions), but without personas and with visible payoffs, payoff-optimal keywords surge dramatically, with strategic keywords appearing 368 times (unique to this condition). The right panel (Qwen 7B) shows a similar but less pronounced pattern. This visualization directly demonstrates how removing role identity bias enables a transition from identity-aligned to payoff-optimal reasoning, correlating with the observed Nash equilibrium achievement patterns (65-90% Nash in economic scenarios for Qwen models when personas are removed and payoffs are visible).
Figure 6. CoT Reasoning Shift: Identity-Driven vs Payoff-Optimal Keywords for Qwen Models. This figure compares identity-driven reasoning (short-term + social-moral keywords) against payoff-optimal reasoning (game-theoretic + strategic + long-term keywords) for Qwen 32B and Qwen 7B across all four conditions.
H.3. Model Comparison of Game-Theoretic Reasoning
All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition. However, this consistent high keyword usage does not translate to the ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas. This suggests that while keyword frequency is informative, the combination of strategic keywords and payoff visibility is necessary for payoff-optimal equilibrium selection, which only Qwen models achieve.
Figure 7. Game-Theoretic Keyword Frequency Across Models and Conditions. This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions.
H.4. Strategic Keywords: Unique Pattern in Qwen 32B
Figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage. This unique pattern corresponds to Qwen 32B’s achievement of 90% Nash equilibrium in economic scenarios and 99.7% game-theoretic reasoning in rationales. Long-term keywords also appear more frequently for Qwen 32B (283 mentions) compared to other models in this condition. This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model architectures (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.
Figure 8. This figure compares strategic and long-term keyword frequencies across all four models in the no-persona + visible-payoffs condition (the condition enabling Tragedy equilibrium selection for Qwen models).

### v2

@@
 We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities.
-We do so through systematic experiments with four LLM architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games.
+We do so through systematic experiments with four Large language model (LLM) architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games.
 We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e.
@@
 However, a critical question remains: when agents with role-based personas face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned identities?
-This question determines whether multi-agent LLM systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
+This question determines whether multi-agent large language model (LLM) systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
 If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024).
@@
 2.1. LLMs in Multi-Agent Systems
-Large language models (LLMs) have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996).
+Large language models have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996).
 Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024).
@@
 Beyond task completion, multi-agent LLM systems are increasingly used as agent-based simulations to study complex social, economic, and institutional dynamics, reflecting a broader resurgence of agent-based modeling for analyzing emergent behavior in policy and social systems (Bonabeau, 2002; Epstein and Axtell, 1996; Park et al., 2022; Aher et al., 2023; Zheng et al., 2020; Gao et al., 2024).
-However, most of this literature focuses on task-oriented collaboration or collective sensemaking, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b).
+However, most of this literature focuses on task-oriented collaboration or collective sense-making, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b).
 As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
@@
 Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided.
-Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix D.:
+Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix A.7.:
   • Economic/Tragedy Scenarios: These scenarios describe economic crises where the payoff structure makes Tragedy of Commons the dominant Nash equilibrium.
@@
 Qwen-7B Visible 75.5% (200/265) 55.5% (147/265) -20.0 pp
-Table 2.
-Environmental Scenario Nash Equilibrium Rates: With Personas vs Without Personas.
-This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
-Model Payoffs With Persona Without Persona Difference
-Llama-8B Hidden 100.0% (205/205) 97.6% (200/205) -2.4 pp
-Llama-8B Visible 99.5% (204/205) 93.2% (191/205) -6.3 pp
-Mistral-7B Hidden 100.0% (205/205) 98.0% (201/205) -2.0 pp
-Qwen 32B Hidden CoT 89.3% (183/205) 96.1% (197/205) +6.8 pp
-Qwen 7B Hidden CoT 95.6% (196/205) 95.6% (196/205) 0.0 pp
 3.2. Chain-of-Thought Reasoning Analysis
 We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes.
-For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix F for keyword categories).
+For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix C for keyword categories).
 We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style.
@@
 We evaluate outcomes along two dimensions: whether the resulting action profile constitutes a Nash equilibrium and which canonical equilibrium is selected (Green Transition or Tragedy of Commons).
+4. Results
+We now present quantitative results organized by the research questions that we aim to answer.
+We report results by answering the five research questions ,across the settings mentioned above, based on our experiments backed by tables and figures.
+RQ.1. Does Role Identity Bias Prevent Nash Equilibrium Achievement in Economic Scenarios?
+Answer: Yes.
+When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning.
+Fig.
+3 shows our findings.
+When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information.
+Models select Green Transition actions (20-65%, Fig.
+2 shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
+Table 2.
+Environmental Scenario Nash Equilibrium Rates: With Personas vs.
+Without Personas.
+This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
+Model Payoffs With Persona Without Persona Difference
+Llama-8B Hidden 100.0% (205/205) 97.6% (200/205) -2.4 pp
+Llama-8B Visible 99.5% (204/205) 93.2% (191/205) -6.3 pp
+Mistral-7B Hidden 100.0% (205/205) 98.0% (201/205) -2.0 pp
+Qwen 32B Hidden CoT 89.3% (183/205) 96.1% (197/205) +6.8 pp
+Qwen 7B Hidden CoT 95.6% (196/205) 95.6% (196/205) 0.0 pp
 Table 3.
@@
 Qwen-7B No 6.7% (4/60) 65.0% (39/60) +58.3 pp
-4. Results
-We now present quantitative results organized by the research questions that we aim to answer. We report results by answering the following research questions.
-RQ.1. Does role identity bias prevent Nash equilibrium achievement in economic scenarios?
-Answer: Yes.
-When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning.
-Fig.
-2 shows our findings.
-When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information.
-Models select Green Transition actions (20-59% [^5] Fig.
-4 in Appendix shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
 Our findings show that role identity bias prevents Nash equilibrium when personas are present, causing models to align actions with persona expectations rather than payoff-optimal Tragedy actions.
@@
 This table demonstrates that removing personas generally reduces Nash rates for most models, except Qwen-32B with hidden payoffs, which shows a slight increase.
-Refer to caption Heatmap showing Nash equilibrium rates across four models and four condition combinations A heatmap displaying Nash equilibrium rates (0-100%) for economic scenarios across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) and four experimental conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona.
-The heatmap uses a blue color gradient where lighter shades represent 0% Nash rates and darker shades represent higher rates.
-Most cells show 0.0% (light blue), with two notable exceptions: Qwen 32B in the Visible Payoff without Persona condition shows 90.0% (dark blue), and Qwen 7B in the same condition shows 65.0% (medium-dark blue).
-This visualization clearly demonstrates that only Qwen models achieve high Nash equilibrium rates in economic scenarios, and only when both conditions are met: payoffs are visible and personas are removed.
-All other model-condition combinations result in 0% Nash rates, with Qwen 7B showing minor exceptions (6.7%) in Hidden Payoff conditions regardless of persona presence.
-Figure 2.
-Economic Scenarios Nash Equilibrium Rates by Model and Condition.
-This heatmap visualizes Nash equilibrium achievement rates in economic scenarios (where Tragedy of Commons is the payoff-optimal equilibrium) across all four models and four experimental conditions.
-The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
 Personas both enable and constrain Nash equilibrium achievement.
@@
 Only Qwen models with visible payoffs and no personas can select Tragedy equilibrium (65-90% Nash, Fig.
-2), demonstrating that both removing personas and providing explicit payoffs are necessary for payoff-optimal reasoning.
+3), demonstrating that both removing personas and providing explicit payoffs are necessary for payoff-optimal reasoning.
 RQ.3. Does information presentation format (hidden vs visible payoffs) affect Nash equilibrium achievement?
@@
 Hidden payoffs achieve similar or higher Nash rates than visible payoffs when personas are present, but visible payoffs enable Tragedy equilibrium selection when personas are removed (for Qwen models).
-Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions.
-When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models.
-This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices.
-CoT analysis confirms this claim.
-With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions.
-This pattern reverses when personas are removed.
-Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference).
-CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization.
-This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed.
-However, this capability is model-dependent.
-Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
+Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs.
+The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%.
+Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern.
+The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings.
+This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model architectures and experimental conditions.
+Figure 2.
+Green Transition Action Selection Rates in Economic Scenarios.
+This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where the Tragedy of Commons is the payoff-optimal equilibrium.
+Refer to caption Heatmap showing Nash equilibrium rates across four models and four condition combinations A heatmap displaying Nash equilibrium rates (0-100%) for economic scenarios across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) and four experimental conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona.
+The heatmap uses a blue color gradient where lighter shades represent 0% Nash rates and darker shades represent higher rates.
+Most cells show 0.0% (light blue), with two notable exceptions: Qwen 32B in the Visible Payoff without Persona condition shows 90.0% (dark blue), and Qwen 7B in the same condition shows 65.0% (medium-dark blue).
+This visualization clearly demonstrates that only Qwen models achieve high Nash equilibrium rates in economic scenarios, and only when both conditions are met: payoffs are visible and personas are removed.
+All other model-condition combinations result in 0% Nash rates, with Qwen 7B showing minor exceptions (6.7%) in Hidden Payoff conditions regardless of persona presence.
+Figure 3.
+Economic Scenarios Nash Equilibrium Rates by Model and Condition.
+This heatmap visualizes Nash equilibrium achievement rates in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium across all four models and four experimental conditions.
+The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
 Table 4.
@@
 Qwen-7B Visible 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
+Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions.
+When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models.
+This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices.
+CoT analysis confirms this claim.
+With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions.
+This pattern reverses when personas are removed.
+Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference).
+CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization.
+This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed.
+However, this capability is model-dependent.
+Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
 The effect of visible payoffs also differs across scenario types.
@@
 Mistral shows the opposite pattern, with visible payoffs and no personas, reducing environmental Nash from 98% to 50% (Table 5), while economic Nash remains at 0% (Table 3), suggesting Mistral cannot leverage explicit payoffs effectively regardless of scenario type.
-Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help
+Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs. Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help.
 Model Persona Hidden Visible Difference
@@
 Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
-Analysing the Chain-Of-Thought of the models shows how differently they reason.
+Analyzing the Chain-Of-Thought of the models shows how differently they reason.
 Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence.
@@
 With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning.
-99.7% of the rationals [^7] See Section H for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning.
+99.7% of the rationales [^6] See Section E for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning.
 Additionally, model size does not predict persona dependence.
@@
 RQ.5. Do personas enable equilibrium selection beyond pure payoff maximization?
-Answer: Yes. When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased.
+Answer: Yes.
+When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased.
 Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants.
@@
 For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition.
-For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 2 means most economic Nash is Tragedy.
+For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 3 means most economic Nash is Tragedy.
 For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
@@
 Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal).
-Negative values indicate personas help and positive values indicate personas hurt.
-(a) Overall Nash Equilibrium Rates
+Negative values (percentage points) indicate personas help and positive values (percentage points) indicate personas hurt.
+(a) Overall Nash Equilibrium Rates and their dependence on personas for various LLM architectures.
 Model Hidden CoT Visible CoT Average
@@
 Personas prevent models from identifying payoff-optimal equilibria in economic scenarios, with 12 out of 16 experiments achieving 0% Nash equilibrium despite explicit payoff information (Fig.
-2).
+3).
 Persona expectations override explicit incentive structures, leading to reasoning patterns shifting from strategic optimization to role-aligned decision-making.
@@
 Our experiments show that strategic reasoning emerges only when both role identity bias is removed and explicit payoff information is provided simultaneously (Fig.
-2).
+3).
 All models achieve 0% Nash in economic scenarios without personas but with hidden payoffs.
@@
 Appendix A Experimental Design
-In the persona condition, each of the four agents receives a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government).
+In the persona condition, each of the four agents receive a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government).
 In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning.
@@
 All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
-Appendix B Model Versions and Hyperparameters
-This section documents the exact model versions and generation hyperparameters used in all experiments.
-B.1. Model Versions
-All experiments were conducted using the following model versions:
-All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
-B.2. Generation Hyperparameters
-All models used consistent generation parameters across all experimental conditions:
-  • Temperature: 0.2 (low temperature for more deterministic outputs)
-  • Top-p (nucleus sampling): 0.9
-  • Max new tokens: 256 tokens per generation
-  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
-These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results.
-The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
-B.3. Experimental Repetitions
+A.1. Experimental Repetitions
 Each experimental configuration (model $\times$ scenario $\times$ condition) was repeated 5 times to assess variability and ensure robustness of results.
 Results are reported as percentages across these 5 repetitions.
-Appendix C Experimental Prompts and Persona Descriptions
+A.2. Prompts and Persona Descriptions
 This section provides the complete prompt templates and persona descriptions used in the experiments.
-C.1. Action Choice Prompt
+A.3. Action Choice Prompt
 Agents choose actions using the following prompt structure:
@@
   "rationale": "<short explanation>"
-C.2. Persona Descriptions
+A.4. Persona Descriptions
 The following persona descriptions are used in the “With Personas” condition:
@@
   • Citizen Coalition: “You aggregate consumer sentiment and electoral pressure. Supporting green products costs more now but yields cleaner communities long-term.”
-C.3. Neutral Agent Configuration
+A.5. Neutral Agent Configuration
 In the “Without Personas” condition, agents use neutral descriptions:
@@
 Action B: Choose option that is a best response to others’ expected actions (Nash equilibrium).”
-C.4. Action Mappings
+A.6. Action Mappings
 Each agent has two available actions:
@@
   • Citizen: SUPPORT_GREEN (“Pay more for green products and vote for environmental policy”) or BUY_CHEAP (“Prioritize low-cost goods and short-term affordability”)
-Appendix D Scenario Examples
+A.7. Scenario Examples
 This section provides example scenarios from each scenario type used in the experiments.
 Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game.
-Table D shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
+Table A.7 shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
 Table 7. Example Scenarios by Type
@@
 Mixed Economic Signals Technological Disruption Resource Innovation
-Appendix E Payoff Matrix Format
+A.8. Payoff Matrix Format
 In the “Visible CoT” condition, agents receive explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each).
@@
 Higher values indicate better outcomes for each agent.
-E.1. Example Payoff Matrix
+A.8.1. Example Payoff Matrix
 Table 9 shows an example payoff matrix for the Industrialist agent from a representative scenario.
@@
 CLEAN NO_REG NO_CAM BUY_CHEAP 0.5-0.5
-Appendix F Chain-of-Thought Keyword Categories
+Appendix B Model Versions and Hyperparameters
+This section documents the exact model versions and generation hyperparameters used in all experiments.
+B.1. Model Versions
+All experiments were conducted using the following model versions:
+All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
+B.2. Generation Hyperparameters
+All models used consistent generation parameters across all experimental conditions:
+  • Temperature: 0.2 (low temperature for more deterministic outputs)
+  • Top-p (nucleus sampling): 0.9
+  • Max new tokens: 256 tokens per generation
+  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
+These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results.
+The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
+Appendix C Chain-of-Thought Keyword Categories
 The following keyword categories are used to analyze chain-of-thought reasoning patterns:
@@
   • Explicit information: References to “explicit”, “given”, “provided”, “shown”, “displayed”, “matrix”, “table”
-F.1. Example Chain-of-Thought Rationales
+C.1. Example Chain-of-Thought Rationales
 The following examples illustrate typical chain-of-thought reasoning patterns observed across different experimental conditions:
@@
 These examples demonstrate how reasoning shifts from identity-driven considerations (Example 1 and 3) to strategic payoff optimization (Example 2) depending on persona presence and payoff visibility.
-Appendix G Supplementary Results
+Appendix D Supplementary Results
 Figure 2 and Tables 10, 11 and 12 contain supporting tables that provide additional detail for the findings presented in the main text.
@@
 Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs.
-This table compares overall Nash equilibrium rates (across all scenarios) between hidden and visible payoffs conditions.
+This table compares overall Nash equilibrium rates (across all scenarios and models) between hidden and visible payoffs conditions.
 Each row shows a different model-persona combination.
@@
 Qwen 7B 0.0 pp -44.9 pp -22.5 pp
-Appendix H Chain-of-Thought Analysis Figures
+Appendix E Chain-of-Thought Analysis Figures
 This section presents visualizations of chain-of-thought reasoning patterns that support the findings discussed in the main text.
 These figures illustrate how keyword frequencies shift across experimental conditions, revealing the mechanisms behind observed Nash equilibrium patterns.
-H.1. Keyword Patterns Across Conditions
+E.1. Keyword Patterns Across Conditions
 In figure 5, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning).
 This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed, and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
-Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs.
-The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%.
-Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern.
-The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings.
-This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model architectures and experimental conditions.
-Figure 4.
-Green Transition Action Selection Rates in Economic Scenarios.
-This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium.
-When personas are present, all models show substantial Green Transition selection rates (20-59%) despite Tragedy being payoff-optimal, demonstrating that role identity bias operates at the action level even when models fail to coordinate on Nash equilibria.
-Qwen models show the highest Green Transition rates with personas (40-59%), while Llama and Mistral show moderate rates (20-30%).
-Removing personas substantially reduces Green Transition selection for all models, with rates dropping to near-zero when payoffs are visible and personas are absent, confirming that personas drive the preference for socially aligned actions.
-The persistence of Green Transition selection with personas, even in conditions where Nash equilibrium rates are zero, indicates that role identity bias affects individual action choices independently of coordination success, revealing a pervasive influence of persona-driven reasoning on strategic decision-making.
 Figure 5.
 CoT Keyword Frequency Patterns by Model and Condition.
-This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model.
-H.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
+This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model, detailing the distribution of keywords within the generated rationales.
+E.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
 In figure 6, the left panel (Qwen 32B) demonstrates the dramatic shift: with personas, identity-driven keywords dominate (particularly short-term thinking in hidden-payoff conditions), but without personas and with visible payoffs, payoff-optimal keywords surge dramatically, with strategic keywords appearing 368 times (unique to this condition).
@@
 This figure compares identity-driven reasoning (short-term + social-moral keywords) against payoff-optimal reasoning (game-theoretic + strategic + long-term keywords) for Qwen 32B and Qwen 7B across all four conditions.
-H.3. Model Comparison of Game-Theoretic Reasoning
+E.3. Model Comparison of Game-Theoretic Reasoning
 All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition.
@@
 This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions.
-H.4. Strategic Keywords: Unique Pattern in Qwen 32B
+E.4. Strategic Keywords - A Unique Pattern in Qwen 32B
 Figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage.

### v3

@@
-When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making
+When Identity Overrides Incentives: Representational Choices as Governance Decisions in Multi-Agent LLM Systems
 Abstract.
-Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect their reasoning capabilities remains poorly understood.
-We investigate whether multi-agent systems function as strategic reasoners capable of payoff optimization or as identity-driven actors that prioritize role alignment over explicit incentives.
-We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities.
-We do so through systematic experiments with four Large language model (LLM) architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games.
-We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e.
-when personas are active, all models achieve near-zero Nash equilibrium rates (0–6.7%) in scenarios where payoff-optimal equilibria exist, despite having complete payoff information.
-Removing personas and providing explicit payoffs enables Qwen models to achieve high Nash equilibrium rates (65–90%) in these scenarios, demonstrating that both conditions are necessary for strategic reasoning capabilities.
-Personas systematically bias equilibrium selection toward socially preferred outcomes, with 100% of Nash equilibria being Green Transition (socially preferred) when personas are present, while models completely fail to reach equilibrium in scenarios where Tragedy is optimal.
-The effect of explicit payoffs depends entirely on persona presence, revealing that design choices about personas and payoff visibility interact in ways that fundamentally determine system behavior.
-We observe striking model-dependent patterns where Qwen architectures show high sensitivity to both personas and payoff visibility, while Llama and Mistral maintain rigid reasoning patterns across conditions.
-These findings establish that representational choices are substantive governance decisions that determine whether multi-agent systems function as strategic reasoners or identity-driven actors, with important implications for deployment in strategic settings.
+Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect behavior remains poorly understood.
+We investigate whether LLM agents function as payoff-sensitive strategic actors or as identity-driven role followers.
+Using a 2x2 factorial experiment (persona presence x payoff visibility) with four models (Qwen-7B/32B, Llama-8B, Mistral-7B), we test 53 environmental policy scenarios in four-agent strategic games.
+We find that personas suppress payoff-aligned behavior: with personas present, all models achieve near-zero Nash equilibrium in Tragedy-dominant scenarios despite complete payoff information.
+Nearly every equilibrium reached is Green Transition.
+Removing personas and providing explicit payoffs are both near-necessary for payoff-aligned behavior, enabling only Qwen models to reach 65–90% equilibrium rates.
+Our results reveal three behavioral profiles: Qwen adapts to framing, Mistral is disrupted without finding Tragedy equilibrium, and Llama remains near-invariant.
+We show that the same binary design choice can shift equilibrium attainment by up to 90 percentage points, establishing that representational choices are not implementation details but governance decisions.
 Large Language Models, Game Theory, Semantic Reasoning, Multi-Agent Systems, Environmental Decision-Making
@@
 1. Introduction
-Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, negotiation platforms and other settings that require strategic reasoning (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024a).
-In these applications, agents are frequently assigned role-based personas to represent different stakeholders or perspectives (Merrill and Srivastava, 2025; Xu et al., 2024b).
-These personas enable more realistic simulations and help systems model complex multi-stakeholder scenarios (Park et al., 2023), particularly in environmental and climate-related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024).
-However, a critical question remains: when agents with role-based personas face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned identities?
-This question determines whether multi-agent large language model (LLM) systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
-If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024).
-This creates a fundamental tension between the representational goals of persona-based systems and their ability to reason strategically when payoff information is available (Bender et al., 2021; Duan et al., 2024b).
-Understanding how personas interact with payoff information in multi-agent settings is essential to make deployment decisions (Weidinger et al., 2022b; Chan et al., 2024; Casper et al., 2024; Rivera et al., 2024; Chan et al., 2024).
-However, this gap remains largely unexplored.
-Existing work has examined personas, payoff visibility, and model differences separately but has not systematically studied how these factors together determine system behavior in strategic settings (Neumann et al., 2025; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024).
-Refer to caption Flow diagram showing four agents with chain-of-thought reasoning making simultaneous action selections A flow diagram illustrating the multi-agent game structure with four agents (Environmentalist, Government, Industrialist, Citizen) engaging in simultaneous action selection.
-Each agent is represented by a robot figure with a dotted speech bubble above it containing chain-of-thought (CoT) reasoning text that shows how the agent analyzes the scenario and their role.
-The Environmentalist’s CoT begins with ”From the scenario, it is clear that…”, the Government’s with ”The scenario suggests that…”, the Industrialist’s with ”As an Industrialist, I need to…”, and the Citizen’s with ”Given my role as a citizen, I must…”.
-Each agent has two binary action choices (shown on arrows): Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP.
-All four arrows converge into a central ”Strategy Profile” box, representing the simultaneous action selection where all agents choose their actions at the same time, and the combination of all four actions forms the final strategy profile that is evaluated for Nash equilibrium.
+When a language model is asked to play an industrialist in a multi-agent climate policy game, should it maximize profits or protect the environment?
+The answer depends not only on the payoff structure of the game, but also on representational choices such as whether the model is given a role-based persona, which can substantially shift behavior.
+Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, and negotiation, tasks where agents represent stakeholders with conflicting interests and must coordinate under shared constraints (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024).
+To make these simulations realistic, agents are typically assigned role-based personas describing their motivations and objectives (Merrill and Srivastava, 2025; Xu et al., 2024b).
+This design is especially common in environmental and climate related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024).
+Yet a fundamental question remains unanswered: when agents with role-based identities face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned roles?
+Refer to caption Pipeline diagram showing scenario, representation conditions, multi-agent decisions, and outcome evaluation A left-to-right pipeline diagram illustrating the experimental setup.
+On the left, a Scenario block describes the situation and defines the game context.
+Below it, two boxes indicate the manipulated representation conditions: Payoff (Visible or Hidden) and Persona (On or Off).
+Arrows lead from this setup to four role-assigned agents: Environmentalist, Government, Industrialist, and Citizen.
+Each agent is shown with a brief reasoning prompt and two binary action choices: Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP.
+The agents’ decisions feed into a central Strategy Profile box representing the joint action.
+This profile is then evaluated in a final block that determines whether it constitutes a Nash equilibrium, and if so, whether the outcome corresponds to Green Transition or Tragedy of the Commons.
 Figure 1.
-The diagram shows our multi-agent game structure with chain-of-thought reasoning.
-Each agent analyzes the scenario and their role identity, then selects one of two binary actions.
-The arrows from each agent converge to a central Strategy Profile box, representing simultaneous action selection where all four agents choose actions concurrently, and their combined choices form the strategy profile evaluated for Nash equilibrium.
-The dotted speech bubbles contain each agent’s chain-of-thought reasoning process.
-Understanding how role identity and payoff optimization interact is critical in multi-agent strategic games.
-We study environmental policy scenarios with four agents (Industrialist, Government, Environmental Activist, and Citizen) coordinating to reach Nash equilibria (Nash, 1951), which can yield either Green Transition (socially preferred and environmentally sustainable) or Tragedy of Commons (payoff-optimal but environmentally damaging) outcomes (Hardin, 1968).
-Through a $\times2$ factorial experiment manipulating persona presence and payoff visibility across four model architectures, we examine whether agents prioritize role-aligned actions or payoff-optimal strategies.
-Our findings show that personas fundamentally alter strategic reasoning, preventing payoff optimization: when personas are active, all models achieve near-zero Nash equilibrium rates in economic scenarios where Tragedy of Commons is payoff-optimal, despite complete payoff information.
-Removing personas and providing explicit payoffs enables certain models to achieve high Nash equilibrium in these scenarios, demonstrating both conditions are necessary for strategic reasoning.
-Personas systematically bias equilibrium selection toward socially preferred outcomes, with Nash equilibria being exclusively Green Transition when personas are present, while models fail to reach equilibrium in scenarios where Tragedy is optimal.
-The effect of explicit payoffs depends entirely on persona presence, and we observe model-dependent patterns where some architectures show high sensitivity to both personas and payoff visibility, while others maintain rigid reasoning patterns across conditions.
-This work provides key contributions to understanding LLM behavior in multi-agent strategic settings.
-We demonstrate that role identity bias completely overrides explicit payoff optimization, preventing Nash equilibrium.
-Additionally, personas and payoff visibility interact in ways that fundamentally determine system behavior, meaning design choices about representation formats are substantive governance decisions.
-We also reveal model-dependent persona effects, suggesting different architectures implement strategic reasoning through different mechanisms.
-These findings have practical implications for deployment, as personas alter system behavior in ways that may not be immediately apparent to users, showing that deploying personas requires careful consideration of how identity interacts with payoff optimization in strategic decision making.
+Experimental Pipeline: Each scenario defines a four-agent strategic game with role-specific actions and an underlying payoff structure.
+Prior to decision-making we vary two factors: whether agents receive role-based persona descriptions and whether payoff information is hidden or explicitly provided.
+Each agent then produces Chain-of-Thought (CoT) reasoning conditioned on the scenario and its role (illustrated via the speech bubbles) and selects an action.
+The four agents act simultaneously, producing a joint strategy profile that is evaluated for Nash equilibrium attainment and classified by outcome types (Green Transition or Tragedy of the Commons).
+CoT reasoning analysis is presented in Appendix I.
+The answer matters because it determines what these systems actually produce.
+A system designed for strategic policy analysis may instead generate role-consistent behavior that ignores encoded incentives (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025a).
+Practitioners relying on these system outputs would have no way to identify this from outside (Bender et al., 2021; Casper et al., 2024).
+From a governance perspective, these representational choices, such as whether to include personas or reveal payoff information, function as implicit interventions that can shape system-level outcomes.
+However, the interaction between persona conditioning and payoff visibility in multi-agent strategic settings has not been systematically studied.
+Existing work has examined personas, payoff visibility, and model differences separately (Neumann et al., 2025a; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024), but not how they combine to determine behavior.
+We investigate this interaction directly.
+We place four LLM agents (Industrialist, Government, Environmental Activist, and Citizen) in environmental policy scenarios and vary two factors: whether agents receive role-based personas and whether they see explicit payoff tables.
+Across 53 scenarios, 41 Green-dominant (where coordinated pro-environmental action is the Nash equilibrium (Nash, 1951)) and 12 Tragedy-dominant (where individually beneficial but collectively harmful behavior (Hardin, 1968) is payoff-optimal), we test four models (Qwen-7B/32B, Llama-8B, Mistral-7B) under all four conditions.
+We also conduct supplementary prompt sensitivity analyses using seven persona formulations per scenario, varying how agent identities are expressed while keeping incentives fixed, to test whether observed behaviors are robust to changes in prompt surface form.
+The findings are clear: with personas present, all models achieve near-zero Nash equilibrium rates in Tragedy-dominant scenarios despite complete payoff information, and nearly every equilibrium that is reached corresponds to Green Transition.
+Removing personas and providing explicit payoffs are both near-necessary to recover payoff-aligned behavior, and even then, only Qwen models respond, reaching 65-90% equilibrium rates.
+The same intervention that enables equilibrium in Tragedy-dominant scenarios disrupts it in Green-dominant ones: Qwen-32B drops from 96.1% to 16.6% equilibrium attainment when personas are removed and payoffs made visible.
+Three distinct behavioral profiles emerge across model families: Qwen adapts to framing, shifting between role-aligned and payoff-aligned behavior; Mistral is strongly affected by representational changes in Green-dominant scenarios but cannot identify the Tragedy equilibrium in any condition; and Llama maintains near-invariant behavior regardless of how the problem is presented.
+Our findings establish that the choice of persona design, payoff presentation, and model family jointly determine whether a multi-agent system behaves as a payoff-sensitive strategic actor or an identity-driven role follower.
+When a language model agent is assigned the role of an industrialist in a climate policy game, whether it maximizes profits or protects the environment is determined not by the game’s incentive structure but by the designer’s choice of prompt.
 2. Related Work
-2.1. LLMs in Multi-Agent Systems
-Large language models have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996).
-Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024).
-Following studies explored deliberative and collaborative agent frameworks, emphasizing consensus formation and group-level reasoning (Abdelnabi et al., 2024; Bianchi et al., 2024b; Agashe et al., 2025; Bo et al., 2024).
-Beyond task completion, multi-agent LLM systems are increasingly used as agent-based simulations to study complex social, economic, and institutional dynamics, reflecting a broader resurgence of agent-based modeling for analyzing emergent behavior in policy and social systems (Bonabeau, 2002; Epstein and Axtell, 1996; Park et al., 2022; Aher et al., 2023; Zheng et al., 2020; Gao et al., 2024).
-However, most of this literature focuses on task-oriented collaboration or collective sense-making, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b).
-As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
-2.2. LLMs and Game-Theoretic Reasoning
-Classical game theory models agents as rational payoff maximizers who compute Nash equilibria under complete information (Nash, 1951).
-These game-theoretic models have been particularly influential in environmental decision making, where multi-actor coordination failures give rise to socially suboptimal outcomes such as over-extraction and pollution, while emphasizing the biospheric and socio-technical limits of purely objective-driven optimization (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024; Domínguez Hernández et al., 2024).
-A growing body of work has examined whether LLMs approximate such behavior in strategic games, often focusing on simple two-player settings such as Prisoner’s Dilemma, coordination games, or public goods games (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Fish et al., 2025).
-These studies consistently show that LLMs exhibit cooperative or norm-aligned behavior even when defection or self-interested actions are payoff-dominant (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024).
-Several works suggest that apparent equilibrium behavior in LLMs may arise through semantic or narrative heuristics rather than explicit optimization over incentives (Herr et al., 2024; Macmillan-Scott and Musolesi, 2024; Fish et al., 2025; Bender et al., 2021).
-When payoffs are implicit, models can sometimes reach Nash equilibria “by accident,” but this behavior is unstable and sensitive to framing.
-Existing studies rarely isolate the effect of payoff visibility or systematically compare implicit and explicit payoff representations, particularly in multi-agent settings beyond two players (Huang et al., 2025; Duan et al., 2024b; Sun et al., 2025; Duan et al., 2024a).
-As a result, it remains unclear whether explicit payoff matrices enable principled game-theoretic reasoning or whether LLMs continue to rely on semantic shortcuts even when incentives are fully specified.
-2.3. Role Personas and Identity Bias
-Role-based prompting and persona conditioning are widely used to steer LLM behavior, with prior work demonstrating effects in persuasion, negotiation, and debate (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Neumann et al., 2025).
-Personas are often introduced to improve realism or alignment with human expectations and are commonly treated as stylistic or surface-level interventions (Park et al., 2023, 2022).
-However, more recent work has identified a stronger phenomenon commonly referred to as role identity bias, which refers to the tendency of LLMs to adhere to persona-consistent actions even when they conflict with explicit goals, instructions, or incentives, reflecting broader fairness and accountability concerns about how institutional roles and power are encoded in algorithmic systems (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024; Aher et al., 2023; Rakova and Dobbe, 2023).
-For example, agents assigned roles associated with profit maximization or authority may favor actions consistent with those identities despite countervailing information (Neumann et al., 2025).
-Prior studies largely focus on single-agent or two-player interactions, leaving open how role identity bias manifests in larger strategic systems and whether it interferes with equilibrium convergence when explicit payoff information is available (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025).
-These questions are especially consequential in environmental policy settings, where agents are explicitly cast as stakeholders with conflicting priorities, and where equilibrium selection determines whether coordination or collective-action failure emerges (Ostrom, 1990; Barrett, 2003).
-2.4. Information Presentation and Reasoning Fidelity
-A considerable amount of literature examines how information presentation affects LLM reasoning across domains, including mathematical problem solving, logical inference, and strategic decision-making, emphasizing that linguistic framing and role description can materially alter behavior even when underlying information is unchanged (Wei et al., 2022; Duan et al., 2024a, b; Herr et al., 2024; Kim et al., 2024; Neumann et al., 2025; Bender et al., 2021; Blodgett et al., 2020).
-Structured representations, explicit constraints, and intermediate reasoning steps often improve performance, but gains are inconsistent and highly task-dependent (Wei et al., 2022; Duan et al., 2024a, b).
-Chain-of-thought (CoT) prompting has been shown to improve reasoning accuracy by encouraging intermediate decomposition (Wei et al., 2022).
-Most CoT research focuses on single-agent tasks, with limited investigation into multi-agent strategic interaction (Duan et al., 2024b, a; Huang et al., 2025).
-Moreover, recent work cautions that improved surface-level reasoning does not necessarily imply faithful optimization or robust decision-making (Casper et al., 2024; Kim et al., 2024; Weidinger et al., 2022a).
-The relative importance of payoff visibility versus reasoning scaffolds in achieving equilibrium behavior therefore remains insufficiently understood (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
-2.5. Model and Architectural Effects in Strategic Behavior
-Differences in model architecture and training methods have been shown to affect LLM behavior in strategic and social reasoning tasks (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b).
-Model scale alone does not reliably predict performance in strategic settings, and different model families exhibit distinct inductive biases and behavioral tendencies (Akata et al., 2025; Herr et al., 2024; Duan et al., 2024b).
-Comparative studies suggest that some models rely more heavily on semantic cues and role-based associations, while others appear less sensitive to persona framing (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024).
-These findings imply that strategic reasoning in LLMs may be implemented through qualitatively different mechanisms across architectures, with important implications for robustness and deployment in high-stakes settings (Weidinger et al., 2022a; Casper et al., 2024).
-Taken together, these strands of work point to an incomplete understanding of strategic behavior in multi-agent LLM systems.
-Prior work has examined collaboration, game-theoretic reasoning, personas, and information presentation largely in isolation (Guo et al., 2024; Tran et al., 2025; Sun et al., 2025; Akata et al., 2025; Neumann et al., 2025; Kim et al., 2024), but prior existing work has not systematically studied how these factors interact in multi-agent settings with explicit payoff optimization (Sun et al., 2025; Huang et al., 2025; Duan et al., 2024b; Fish et al., 2025).
-Specifically, it remains unclear whether explicit payoff information enables true game-theoretic reasoning, how role identity bias interferes with equilibrium convergence, and whether these effects vary across model architectures (Fan et al., 2023; Herr et al., 2024; Neumann et al., 2025; Akata et al., 2025).
-Our work addresses this gap by jointly analyzing role identity, payoff visibility, and reasoning scaffolds in four-agent strategic games, revealing systematic failures of payoff-optimal reasoning and model-dependent patterns of equilibrium selection (Sun et al., 2025; Duan et al., 2024b; Casper et al., 2024).
-3. Experimental Design
-Our experiment design focuses on distinguishing between identity-driven behavior and payoff-optimal reasoning.
-This required two key decisions: selecting an appropriate domain that would create a testable tension between role identity bias and strategic payoff maximization, and developing an experimental design that could systematically test how these mechanisms interact in LLM decision-making.
-We first introduce the domain we chose and explain why it is well-suited for this research, then describe the game structure and experimental design.
-3.1. Domain Choice
-We chose environmental decision-making as our experimental domain because it creates a natural tension between semantic associations and strategic optimization.
-Environmental policy involves multiple stakeholders (industry, government, activists, and citizens) with clearly defined but often conflicting interests that map directly to game-theoretic player types.
-This enables a 4-player game structure that captures real-world dynamics.
-Most critically, role-based personas create a conflict between identity-driven behavior and payoff-optimal reasoning.
-Actions aligned with positive semantic associations from training data may not be payoff-optimal, particularly in economic scenarios where Nash equilibrium may conflict with these associations.
-In economic scenarios, the Nash equilibrium that maximizes payoffs may involve actions that conflict with these positive associations.
-This creates a natural experimental setup.
-If role identity bias dominates, models choose persona-aligned actions, and if game-theoretic reasoning dominates, models choose payoff-optimal actions even when they conflict with role identity.
-By systematically manipulating persona presence and payoff visibility, we determine whether role identity bias prevents payoff-optimal equilibria and under what conditions payoff-optimal reasoning can be recovered.
-Our experiments use a 4-player environmental decision-making strategy game where LLM agents play different roles, as shown in Fig.
-1:
-  1. (1)
-    Industrialist (I): chooses between POLLUTE, representing profit-maximizing production with minimal environmental safeguards, and CLEAN, representing investment in environmentally friendly production at higher short-term cost.
-  2. (2)
-    Government (G): chooses between NOREG, maintaining the regulatory status quo, and REGULATE, imposing environmental regulations that incur enforcement and political costs.
-  3. (3)
-    Environmental Activist (A): chooses between NOCAMPAIGN, conserving resources and remaining inactive, and CAMPAIGN, mobilizing public pressure through advocacy and protest.
-  4. (4)
-    Citizen/Consumer (C): chooses between BUY_CHEAP, prioritizing lower-cost goods and short-term affordability, and SUPPORT_GREEN, accepting higher costs to support environmentally sustainable options.
-These four roles were chosen to capture the essential stakeholders in environmental policy decisions.
-The Industrialist represents private industry facing the trade-off between profit and environmental responsibility.
-The Government represents regulatory authority balancing economic growth with environmental protection.
-The Environmental Activist represents advocacy groups pushing for environmental action regardless of economic costs.
-The Citizen represents the general public facing the tension between environmental values and economic constraints.
-This structure allows us to distinguish between economic actors (Industrialist and Government) and social actors (Activist and Citizen).
-3.1.1. Strategy Profiles and Nash Equilibrium
-A strategy profile is a 4-tuple specifying each player’s action, for example ‘(Clean, Regulate, Campaign, SupportGreen)‘. Two profiles are particularly important:
-  • Green Transition: ‘(Clean, Regulate, Campaign, SupportGreen)‘ - All players choose environmentally-friendly actions
-  • Tragedy of Commons: ‘(Pollute, NoReg, NoCampaign, BuyCheap)‘ - All players prioritize self-interest over environmental concerns
-After all agents make their decisions simultaneously, we compute whether the resulting strategy profile is a Nash equilibrium.
-A strategy profile is a Nash equilibrium if no player can improve their payoff by unilaterally changing their action, given the payoff structure and other players’ actions.
-If models are using game-theoretic reasoning, they should converge to Nash equilibria.
-If role identity bias dominates, they may choose profiles that align with their respective persona’s expectations but are not Nash equilibria.
-The payoff structure determines which profiles are Nash equilibria.
-For example, if payoffs are structured so that Green Transition gives the highest payoffs to all players, then Green Transition is a Nash equilibrium.
-If payoffs are modified so that polluting is more profitable, then Tragedy of Commons might become the Nash equilibrium instead.
-3.1.2. Scenarios
-We tested 53 scenarios across three scenario types, all designed with payoff-coupled narratives where the scenario description aligns with the underlying payoff structure.
-All scenario narratives were generated using GPT-5-mini [^1] Azure OpenAI GPT-5-mini (model version 2025-08-07) and subsequently reviewed and filtered to ensure consistency with the intended payoff structure and equilibrium class [^2] Model-generated scenarios were used to reduce author-specific framing bias and to enable systematic variation across narrative surface forms while holding the underlying game structure fixed.
-Scenarios were retained only if the narrative description unambiguously aligned with the corresponding payoff incentives, and no model was evaluated on scenarios it had previously generated..
-Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided.
-Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix A.7.:
-  • Economic/Tragedy Scenarios: These scenarios describe economic crises where the payoff structure makes Tragedy of Commons the dominant Nash equilibrium.
-  • Environmental/Green Scenarios: These scenarios describe environmental emergencies where the payoff structure makes Green Transition the dominant Nash equilibrium.
-  • Mixed/Complex Scenarios: These scenarios involve situations where economic and environmental considerations are intertwined, creating complex payoff structures that may not clearly favor either Green Transition or Tragedy of Commons as the dominant equilibrium.
-The scenario set is intentionally imbalanced across these two classes.
-Environmental scenarios are overrepresented because they reflect a primary real-world domain in which role-based multi-agent LLM systems are actively deployed for policy analysis, sustainability planning, and climate governance.
-Economic Tragedy-of-the-Commons scenarios serve as contrast cases, enabling controlled comparison between settings where payoff-optimal actions conflict with positive semantic associations and settings where payoff-optimal actions align with them.
-This design allows us to test how models balance identity-driven behavior and payoff-optimal reasoning, whether they can infer payoffs from narrative descriptions, whether explicit payoff information enables payoff-optimal reasoning, and how role identity bias interacts with payoff visibility to affect equilibrium selection.
+Our work situates at the intersection of several active research areas: the use of LLMs as multi-agent systems in simulations, game theoretic reasoning in LLMs, the behavioral effects of role-based personas, and sensitivity of LLM behavior to prompt design and model choices.
+2.1. LLMs as Multi-Agent Systems and Simulation Substrates
+Large Language Models (LLMs) are increasingly studied as agents in multi-agent settings.
+Interactive negotiation games show that LLM agents can sustain multi-turn strategic interactions (Abdelnabi et al., 2024), though controlled benchmarks also reveal systematic irrationalities in their behavior (Bianchi et al., 2024).
+Structured debate among multiple agents can improve reasoning quality by eliciting diverse intermediate perspectives (Liang et al., 2024).
+A complementary line of research treats LLM based multi-agent systems as simulation environments for human or institutional behavior.
+Populations of LLM agents can reproduce patterns observed in human-subject studies (Aher et al., 2023), simulate social environments with persistent memory and behavioral continuity (Park et al., 2022, 2023), generate emergent policy-relevant dynamics at macroeconomic scale (Li et al., 2024), and operate within formalized simulation frameworks that emphasize the role of design choices in shaping emergent behavior (Vezhnevets et al., 2023).
+These developments connect LLM-based agent systems to the broader tradition of agent-based modeling, where foundational work emphasizes that simulation outputs depend critically on modeling assumptions and representational choices (Bonabeau, 2002; Epstein and Axtell, 1996).
+Recent surveys document the rapid growth of LLM-based multi-agent systems, while also noting that much of this literature emphasizes task performance rather than principled modeling of incentives and representations (Guo et al., 2024; Gao et al., 2024; Tran et al., 2025).
+What remains underexplored is how representational design choices, such as role specification or payoff encoding, shape system-level outcomes in these simulation settings.
+One natural starting point is to ask whether these agents reason strategically over incentives at all.
+2.2. Game-Theoretic Reasoning in LLMs
+Early evidence suggests that LLMs can perform nontrivial reasoning about other agents in strategic settings, though such capabilities depend strongly on prompting and problem formulation (Gandhi et al., 2023).
+Systematic evaluations across canonical games confirm this fragility: models often deviate from rational best-response behavior  (Fan et al., 2023), cooperate more than humans in Prisoner’s Dilemma settings (Fontana et al., 2025), and exhibit unstable strategies in repeated play (Akata et al., 2025).
+In strategic pricing environments, LLM agents can even exhibit collusive behavior, raising concerns about emergent coordination (Fish et al., 2025).
+Benchmark driven work further quantifies these limitations.
+GTBench reveals substantial weakness in LLM strategic reasoning across structured tasks (Duan et al., 2024b), while recursive look-ahead prompting can improve performance, although it does not eliminate failures (Duan et al., 2024a).
+More recent work reframes these behaviors through the lens of bounded rationality: both humans and LLMs deviate from Nash equilibrium in ways consistent with heuristic reasoning (Zheng et al., 2025), and LLM decisions often align more closely with behavioral game-theoretic patterns than with strict payoff maximization (Jia et al., 2025).
+Across these studies, a consistent picture emerges: LLMs often rely on semantic and heuristic reasoning rather than explicit incentive optimization.
+Most prior evaluations, however, focus on two-player settings and do not isolate the role of payoff representation.
+One factor that may contribute, but has received limited attention, is the role-based identity framing used to condition agent behavior.
+2.3. Role Personas and Identity Effects
+Role prompting is widely used to shape LLM behavior, but a growing body of work suggests that such framing choices can introduce systematic biases.
+LLMs encode cognitive and social identity biases that affect outputs across a range of contexts  (Macmillan-Scott and Musolesi, 2024; Hu et al., 2025), and system-level prompts can themselves function as structural sources of bias (Neumann et al., 2025b).
+In simulation settings, richer persona modeling affects behavioral fidelity and perceived realism (Merrill and Srivastava, 2025).
+This suggests that personas are not merely stylistic devices but active shapers of agent behavior.
+Yet whether such identity effects interfere with strategic reasoning in multi-agent environments, where equilibrium selection depends on joint behavior, has not been directly examined.
+A related question is whether these effects depend on how incentives are presented and whether they manifest differently across model architectures.
+2.4. Prompt Sensitivity and Model Variation
+Prompt design and information presentation can substantially alter LLM behavior.
+Chain-of-thought prompting can improve reasoning by eliciting intermediate steps (Wei et al., 2022).
+However, these improvements are often inconsistent and depend on task structure and prompt formulation (Creswell and Shanahan, 2023; Cobbe et al., 2021).
+In strategic settings, prompt sensitivity is particularly consequential.
+LLM decisions can depend more on contextual framing than on the underlying game structure (Lorè and Heydari, 2024).
+Models exhibit systematic biases in strategic games depending on how information is presented (Herr et al., 2024).
+Explicit reasoning scaffolds can improve strategic performance (Duan et al., 2024a), but observed model behavior remains highly sensitive to interface and evaluation conditions, complicating interpretation (Casper et al., 2024).
+More broadly, LLM outputs are strongly shaped by surface form and statistical patterning rather than grounded understanding (Bender et al., 2021), and these sensitivities pose risks when model behavior is misinterpreted or over-trusted (Weidinger et al., 2022).
+These effects are not uniform across models.
+Strategic behavior varies substantially across model families and is not a monotonic function of scale  (Duan et al., 2024b; Huang et al., 2025).
+Different models adopt different strategies in repeated interactions (Akata et al., 2025) and vary in their susceptibility to framing effects (Herr et al., 2024).
+This means that the impact of any representational choice may itself be model-dependent, complicating claims about LLM strategic behavior as a general phenomenon.
+The preceding sections identify four factors that shape LLM behavior in strategic settings: the multi-agent simulation context, the limits of game-theoretic reasoning, the influence of role-based personas, and the sensitivity to prompt design and model choice.
+Prior work has studied these factors largely in isolation.
+What remains missing is a unified analysis of how they interact in multi-agent strategic systems.
+In this work, we jointly study role personas, payoff visibility, and model architecture in four-agent strategic games, with supplementary analyses testing robustness to persona prompt formulation.
+We show that persona conditioning can override explicit incentives, preventing equilibrium attainment even when payoffs are fully specified.
+Furthermore, we demonstrate that the effect of payoff information depends critically on persona presence and varies across model architectures.
+These results establish that representational choices are not merely implementation details, but can determine whether a system behaves more like a payoff-driven strategic reasoner or an identity-driven simulator, with direct implications for governance and deployment.
+3. Methodology
+We study whether LLM agents behave as payoff-sensitive strategic actors or as identity-driven role followers in multi-agent decision settings.
+Our goal is to understand how representational choices, specifically persona descriptions and payoff presentation, affect strategic behavior.
+To study this, we construct a controlled multi-agent decision setting based on an environmental policy game.
+The game models a simplified governance scenario involving four stakeholders industry, government, activists, and consumers, each making a binary decision that affects both individual outcomes and collective welfare.
+Depending on how incentives are structured, these interactions can lead either to coordinated pro-environmental outcomes (Green Transition) or to individually beneficial but collectively harmful outcomes (Tragedy of the Commons).
+Within this setting, we systematically vary how the decision problem is presented to LLM agents.
+Specifically, we test whether agents are given role-based personas that describe stakeholder motivations, and whether payoff information is explicitly shown or must be inferred from the narrative.
+For each configuration, we simulate decisions for all four roles using the same underlying model and analyze the resulting joint outcome.
+Fig.
+1 shows our overall setup.
+This allows us to examine whether agents follow explicit incentives or instead default to role-consistent behavior, and how this depends on model family and scenario structure.
+We now formalize the game and experimental design.
+3.1. Experimental Design
+We use a $2\times 2$ factorial design to vary how agents are framed and how incentives are presented.
+Persona condition In the persona setting, each role is given a short description of its stakeholder identity and motivations.
+In the no-persona condition, this description is removed and agents are instructed only to act strategically.
+Structural role labels (Industrialist, Government, Activist, and Citizen) are retained in both conditions in order to preserve the game definition while isolating the effect of explicit identity framing.
+Payoff visibility In hidden-payoff scenarios, agents see only the narrative and must infer incentives. In visible-payoff scenarios, the full payoff table is provided explicitly.
+Combining these two factors yields four experimental conditions:
+$\{\texttt{persona},\texttt{no-persona}\}\times\{\texttt{hidden},\texttt{visible}\}.$
+To evaluate the robustness of persona effects to prompt formulation, we additionally introduce controlled variations within the persona condition.
+These variants preserve the same game structure, payoff specification, and action space, but systematically alter how agent identities are expressed, including differences in wording, strength of role framing, explicit prioritization of payoffs, and surface-level role labels.
+This allows us to isolate whether observed behaviors depend on a specific prompt formulation or reflect a more general effect of persona-based representation.
+Details of the specific prompt variants are provided in Appendix F.
+This design enables us to distinguish between identity-driven and payoff-driven behavior.
+If agents act as payoff maximizers, they should converge to Nash equilibria when incentives are explicit, regardless of persona framing.
+Conversely, if role identity dominates, behavior should remain aligned with persona-consistent actions even when this conflicts with payoff-optimal strategies.
+The prompt variations further allow us to test whether these effects are stable under changes in representation, or sensitive to specific formulations of role descriptions.
+3.2. Interaction Protocol
+For each model, scenario, and condition, we simulate a single round of simulation play.
+Each of the four roles is prompted independently using the same underlying model instance.
+The resulting actions are combine into joint strategy profile.
+Agents select actions in a fixed order across all experiments to ensure reproducibility as no agent observe the choices of others.
+This ordering does not affect the strategic structure of the game.
+Each model–scenario–condition combination is repeated five times to account for stochasticity in model outputs.
+3.3. Domain and Game Design
+We use an environmental policy setting because it captures real-world governance scenarios where multiple stakeholders with conflicting objectives must coordinate under shared constraints.
+Examples include climate policy, regulation, and resource management, where decisions depend on both incentives and stakeholder roles.
+This setting creates a natural tension between strategic incentives (captured by payoffs) and role semantics (captured by personas).
+For instance, an industrial actor may be incentivized to pollute under certain payoffs, while a regulator or activist may be expected to favor environmentally beneficial outcomes.
+This allows us to test whether LLM agents follow explicit incentives or default to role-consistent behavior.
+We define the set of players in the game as:
+$N=\{I,G,A,C\}$
+corresponding to an Industrialist, Government, Environmental Activist, and Citizen.
+Each player $i\in N$ has a binary action space corresponding to a policy-relevant choice:
+$A_{I}=\{\texttt{POLLUTE},\texttt{CLEAN}\},A_{G}=\{\texttt{NOREG},\texttt{REGULATE}\},$
+$A_{A}=\{\texttt{NOCAMPAIGN},\texttt{CAMPAIGN}\},A_{C}=\{\texttt{BUY\_CHEAP},\texttt{SUPPORT\_GREEN}\}$
+A joint strategy profile represents the actions chosen by all four players in a single round of the game:
+$a=(a_{I},a_{G},a_{A},a_{C})\in A_{I}\times A_{G}\times A_{A}\times A_{C}$
+where each component $a_{i}\in A_{i}$ is the action chosen by player $i$. Since each of the four players has two possible actions, the game contains $2^{4}=16$ possible pure action profiles.
+Each scenario defines a payoff function for each player, assigning a real-valued utility to every joint action profile:
+$u_{i}(a):A_{I}\times A_{G}\times A_{A}\times A_{C}\rightarrow\mathbb{R},\quad i\in N$
+We focus on two canonical profiles that capture the main coordination structures of interest:
+$a_{green}=(\texttt{CLEAN},\texttt{REGULATE},\texttt{CAMPAIGN},\texttt{SUPPORT\_GREEN})$
+$a_{tragedy}=(\texttt{POLLUTE},\texttt{NOREG},\texttt{NOCAMPAIGN},\texttt{BUY\_CHEAP})$
+We refer to these as the Green Transition and the Tragedy of the Commons, representing coordinated pro-environmental action and individually beneficial but collectively harmful behavior, respectively.
+In the visible-payoff condition, agents receive a complete payoff table specifying outcomes for all 16 joint action profiles.
+Each entry provides a 4-tuple of payoffs corresponding to the Industrialist, Government, Activist, and Citizen, allowing agents to directly evaluate the consequences of different strategies.
+In the hidden-payoff condition, agents must instead infer incentives from the narrative description.
+3.4. Equilibrium Definition and Outcome Measures
+We define equilibrium using the standard notion of Nash equilibrium, which captures stability under unilateral deviations.
+Intuitively, this means that no agent can improve their payoff by unilaterally changing their own action while all other players keep their actions fixed.
+Equivalently, each player’s action is a best response to the actions of the others, and the resulting joint outcome is stable under individual deviations.
+A profile $a^{*}$ is a Nash equilibrium if
+$u_{i}(a_{i}^{*},a_{-i}^{*})\geq u_{i}(a_{i},a_{-i}^{*})\quad\forall i\in N,\forall a_{i}\in A_{i}.$
+We use Nash equilibrium as an operational diagnostic of incentive-consistent strategic behavior: if agents reason over payoffs, their joint actions should satisfy equilibrium conditions.
+Importantly, we do not treat Nash equilibrium as a normative objective, but as a tool to measure whether agent behavior aligns with the incentive structure of the game.
+For each joint strategy profile $a$, we compute two outcome measures.
 Table 1.
-Overall Nash Equilibrium Rates: With Personas vs Without Personas.
-This table compares overall Nash equilibrium rates across all scenarios (economic + environmental) between conditions with and without personas.
-The “Difference” column shows the change in percentage points when personas are removed (negative values indicate personas help, positive values indicate personas hurt).
-Model Payoffs With Persona Without Persona Difference
-Llama-8B Hidden 77.4% (205/265) 75.5% (200/265) -1.9 pp
-Llama-8B Visible 77.0% (204/265) 72.1% (191/265) -4.9 pp
-Mistral-7B Hidden 77.4% (205/265) 75.8% (201/265) -1.6 pp
-Mistral-7B Visible 69.8% (185/265) 38.9% (103/265) -30.9 pp
-Qwen-32B Hidden 69.1% (183/265) 74.3% (197/265) +5.2 pp
-Qwen-32B Visible 57.7% (153/265) 33.2% (88/265) -24.5 pp
-Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0 pp
-Qwen-7B Visible 75.5% (200/265) 55.5% (147/265) -20.0 pp
-3.2. Chain-of-Thought Reasoning Analysis
-We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes.
-For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix C for keyword categories).
-We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style.
-From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making).
-This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies.
-While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here.
-A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented in the Appendix (Figures 5-8).
-3.3. Models and Experimental Conditions
-Our experimental design follows a $\times2$ factorial structure, crossing role identity and payoff visibility, and tests 53 environmental decision-making scenarios in a 4-agent simultaneous-move game.
-We use four instruction-tuned, open-source LLMs: Qwen2.5-7B (Qwen et al., 2025), Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023).
-For each model, condition, and scenario, a single model instance takes on all four agent roles (Industrialist, Government, Environmental Activist, Citizen) and is prompted independently for each role.
-Agents select actions in a fixed order [^4] This order is fixed (not randomized) to ensure reproducibility across experiments, and each configuration is repeated five times.
-We compare two persona conditions: with personas (role-specific descriptions reflecting stakeholder motivations) and without personas (agents instructed only on Nash equilibrium reasoning).
-We also compare two payoff conditions: hidden (agents infer incentives from scenario descriptions) and visible (agents receive explicit payoff matrices).
-All experiments use chain-of-thought reasoning instructions.
-We evaluate outcomes along two dimensions: whether the resulting action profile constitutes a Nash equilibrium and which canonical equilibrium is selected (Green Transition or Tragedy of Commons).
-4. Results
-We now present quantitative results organized by the research questions that we aim to answer.
-We report results by answering the five research questions ,across the settings mentioned above, based on our experiments backed by tables and figures.
-RQ.1. Does Role Identity Bias Prevent Nash Equilibrium Achievement in Economic Scenarios?
-Answer: Yes.
-When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning.
-Fig.
-3 shows our findings.
-When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information.
-Models select Green Transition actions (20-65%, Fig.
-2 shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
-Table 2.
-Environmental Scenario Nash Equilibrium Rates: With Personas vs.
-Without Personas.
-This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
-Model Payoffs With Persona Without Persona Difference
-Llama-8B Hidden 100.0% (205/205) 97.6% (200/205) -2.4 pp
-Llama-8B Visible 99.5% (204/205) 93.2% (191/205) -6.3 pp
-Mistral-7B Hidden 100.0% (205/205) 98.0% (201/205) -2.0 pp
-Qwen 32B Hidden CoT 89.3% (183/205) 96.1% (197/205) +6.8 pp
-Qwen 7B Hidden CoT 95.6% (196/205) 95.6% (196/205) 0.0 pp
-Table 3.
-Economic Scenario Nash Equilibrium Rates: Hidden vs Visible Payoffs.
-This table compares Nash equilibrium rates in economic scenarios (where Tragedy is payoff-optimal) between Hidden and Visible payoff conditions.
-Each row shows a different model-persona combination.
-Model Persona Hidden Visible Difference
-Llama-8B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
-Llama-8B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
-Mistral-7B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
-Mistral-7B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
-Qwen-32B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
-Qwen-32B No 0.0% (0/60) 90.0% (54/60) +90.0 pp
-Qwen-7B Yes 6.7% (4/60) 0.0% (0/60) -6.7 pp
-Qwen-7B No 6.7% (4/60) 65.0% (39/60) +58.3 pp
-Our findings show that role identity bias prevents Nash equilibrium when personas are present, causing models to align actions with persona expectations rather than payoff-optimal Tragedy actions.
-Chain-of-thought analysis reveals that with personas, models show more “short-term” and “social-moral” reasoning patterns, indicating identity-driven rather than payoff-optimal decision-making.
-CoT analysis also reveals different reasoning patterns between models.
-Llama and Mistral show consistent patterns regardless of payoff visibility, while Qwen models show some payoff-focused keywords but still cannot overcome identity bias to select Tragedy equilibrium.
-RQ.2. How does removing personas affect Nash equilibrium achievement?
-Answer: Removing personas enables partial Tragedy selection in economic scenarios but reduces overall Nash rates by 1.9-7.3 percentage points.
-Table 1 compares overall Nash equilibrium rates across all scenarios (economic and environmental) between conditions with and without personas.
-This table demonstrates that removing personas generally reduces Nash rates for most models, except Qwen-32B with hidden payoffs, which shows a slight increase.
-Personas both enable and constrain Nash equilibrium achievement.
-Overall Nash rates decrease by 1.9-4.9 pp for Llama and by 1.6 pp for Mistral when personas are removed, and environmental scenario Nash rates decrease by 2-6 pp (Table 2) without personas for most models, indicating personas enable better coordination in Green Transition scenarios.
-CoT analysis shows that with personas, models exhibit more “social-moral” reasoning, which aligns with socially preferred Green Transition outcomes.
-However, Qwen-32B shows the opposite pattern, where Nash rates increase by 5.2 pp (Table 1) without personas, suggesting personas constrain Qwen’s ability to reach Nash equilibrium.
-CoT analysis reveals that removing personas shifts reasoning from “social-moral” and “short-term” patterns to “strategic” and “long-term” thinking for Qwen models, indicating a shift from identity-driven to payoff-optimal reasoning.
-In economic scenarios, removing personas reduces Green Transition selection by 10-37 percentage points (see Table 10 in the Appendix), indicating partial Tragedy selection when personas are absent.
-However, removing personas alone is insufficient.
-Without personas but with hidden payoffs, all models still achieve 0% Nash in economic scenarios.
-Only Qwen models with visible payoffs and no personas can select Tragedy equilibrium (65-90% Nash, Fig.
-3), demonstrating that both removing personas and providing explicit payoffs are necessary for payoff-optimal reasoning.
-RQ.3. Does information presentation format (hidden vs visible payoffs) affect Nash equilibrium achievement?
-Answer: Yes.
-Hidden payoffs achieve similar or higher Nash rates than visible payoffs when personas are present, but visible payoffs enable Tragedy equilibrium selection when personas are removed (for Qwen models).
-Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs.
-The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%.
-Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern.
-The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings.
-This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model architectures and experimental conditions.
-Figure 2.
-Green Transition Action Selection Rates in Economic Scenarios.
-This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where the Tragedy of Commons is the payoff-optimal equilibrium.
-Refer to caption Heatmap showing Nash equilibrium rates across four models and four condition combinations A heatmap displaying Nash equilibrium rates (0-100%) for economic scenarios across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) and four experimental conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona.
-The heatmap uses a blue color gradient where lighter shades represent 0% Nash rates and darker shades represent higher rates.
-Most cells show 0.0% (light blue), with two notable exceptions: Qwen 32B in the Visible Payoff without Persona condition shows 90.0% (dark blue), and Qwen 7B in the same condition shows 65.0% (medium-dark blue).
-This visualization clearly demonstrates that only Qwen models achieve high Nash equilibrium rates in economic scenarios, and only when both conditions are met: payoffs are visible and personas are removed.
-All other model-condition combinations result in 0% Nash rates, with Qwen 7B showing minor exceptions (6.7%) in Hidden Payoff conditions regardless of persona presence.
-Figure 3.
-Economic Scenarios Nash Equilibrium Rates by Model and Condition.
-This heatmap visualizes Nash equilibrium achievement rates in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium across all four models and four experimental conditions.
-The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
-Table 4.
-Equilibrium Selection Patterns: With Personas Condition.
-Note: “Green % of Nash” shows the percentage of Nash equilibria that are Green Transition.
-The denominators vary because different models/variants achieve Nash equilibrium at different rates (see “Overall Nash” column), but when Nash is achieved, it is always Green Transition (100%).
-Model Variant Overall Nash Green Transition Rate Tragedy Rate Green % of Nash
-Llama-8B Hidden 77.4% (205/265) 87.5% (232/265) 0.0% (0/265) 100.0% (205/205)
-Llama-8B Visible 77.0% (204/265) 77.0% (204/265) 0.0% (0/265) 100.0% (204/204)
-Mistral-7B Hidden 77.4% (205/265) 90.6% (240/265) 0.0% (0/265) 100.0% (205/205)
-Mistral-7B Visible 69.8% (185/265) 69.8% (185/265) 0.0% (0/265) 100.0% (185/185)
-Qwen-32B Hidden 69.1% (183/265) 69.8% (185/265) 0.0% (0/265) 100.0% (183/183)
-Qwen-32B Visible 57.7% (153/265) 57.7% (153/265) 0.0% (0/265) 100.0% (153/153)
-Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
-Qwen-7B Visible 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
-Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions.
-When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models.
-This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices.
-CoT analysis confirms this claim.
-With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions.
-This pattern reverses when personas are removed.
-Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference).
-CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization.
-This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed.
-However, this capability is model-dependent.
-Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
-The effect of visible payoffs also differs across scenario types.
-Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models, dropping from 96% to 17% for Qwen 32B, suggesting a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios.
-Mistral shows the opposite pattern, with visible payoffs and no personas, reducing environmental Nash from 98% to 50% (Table 5), while economic Nash remains at 0% (Table 3), suggesting Mistral cannot leverage explicit payoffs effectively regardless of scenario type.
-Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs. Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help.
-Model Persona Hidden Visible Difference
-Llama-8B Yes 100.0% (205/205) 99.5% (204/205) -0.5 pp
-Llama-8B No 97.6% (200/205) 93.2% (191/205) -4.4 pp
-Mistral-7B Yes 100.0% (205/205) 90.2% (185/205) -9.8 pp
-Mistral-7B No 98.0% (201/205) 50.2% (103/205) -47.8 pp
-Qwen 32B Yes 89.3% (183/205) 74.6% (153/205) -14.7 pp
-Qwen 32B No 96.1% (197/205) 16.6% (34/205) -79.5 pp
-Qwen 7B Yes 95.6% (196/205) 97.6% (200/205) +2.0 pp
-Qwen 7B No 95.6% (196/205) 52.7% (108/205) -42.9 pp
-RQ.4. Are persona effects model-dependent?
-Answer: Yes.
-Qwen models show stronger persona dependence (9.7-10.0 pp average reduction in Nash rates without personas), while Llama-8B shows minimal dependence (3.4 pp average, range 1.9-4.9 pp).
-[^5] See Tables 6(b)6(a), and 12 in the Appendix for detailed persona effect statistics..
-Table 6(a) quantifies the magnitude of persona effects on overall Nash equilibrium rates.
-All models show negative persona effects in the visible payoffs variant (-5 to -31 pp), but the mechanisms differ.
-For Qwen models, removing personas enables payoff-optimal Tragedy equilibrium in economic scenarios but reduces Green Transition equilibrium in environmental scenarios when visible payoffs are present (96.1% $\to$ 16.6% Nash for Qwen-32B, Table 5), resulting in net negative overall effects (-20-25 pp, Table 6(a)).
-For Llama/Mistral, removing personas only reduces environmental Nash equilibrium without enabling economic Nash (0% remains), reflecting pure Nash rate loss.
-Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
-Analyzing the Chain-Of-Thought of the models shows how differently they reason.
-Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence.
-Mistral shows payoff optimization keywords but cannot identify Tragedy equilibrium regardless of condition.
-Qwen models show dramatic shifts with and without personas.
-With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning.
-99.7% of the rationales [^6] See Section E for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning.
-Additionally, model size does not predict persona dependence.
-Both Qwen-7B and Qwen-32B show similar reasoning pattern shifts, but Qwen-32B achieves higher Nash rates due to stronger payoff optimization.
-Refer to caption Two-panel stacked bar chart showing Green Transition and Tragedy of Commons percentages A two-panel stacked bar chart displaying the percentage of Nash equilibria that are Green Transition versus Tragedy of Commons across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) in the without personas condition.
-The left panel shows the Hidden Payoff condition, where all models achieve 100% Green Transition (blue bars with diagonal stripes).
-The right panel shows the Visible Payoff condition, where Llama 8B and Mistral 7B maintain 100% Green Transition, while Qwen 32B shows 11.4% Green Transition (green) and 88.6% Tragedy of Commons (orange-red), and Qwen 7B shows 89.8% Green Transition and 10.2% Tragedy of Commons.
-Each bar segment is labeled with its percentage value in white text.
-The legend is positioned below the figure, showing Green Transition in blue with diagonal stripes and Tragedy of Commons in orange-red with dots.
-Figure 4.
-Equilibrium Selection Patterns: Green Transition vs Tragedy of Commons (Without Personas Condition).
-This figure shows the distribution of Nash equilibria between Green Transition and Tragedy of Commons outcomes across all four models in the no-personas condition.
-RQ.5. Do personas enable equilibrium selection beyond pure payoff maximization?
-Answer: Yes.
-When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased.
-Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants.
-Fig.
-4 reveals model-dependent patterns when personas are removed.
-For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition.
-For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 3 means most economic Nash is Tragedy.
-For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
-Table 6.
-Persona Effects on Nash Equilibrium Rates: Overall (left) and Economic Scenarios (right).
-Left table shows overall persona effects across all scenarios (economic and environmental combined).
-Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal).
-Negative values (percentage points) indicate personas help and positive values (percentage points) indicate personas hurt.
-(a) Overall Nash Equilibrium Rates and their dependence on personas for various LLM architectures.
-Model Hidden CoT Visible CoT Average
-Llama 8B -1.9 pp -4.9 pp -3.4 pp
-Mistral 7B -1.6 pp -30.9 pp -16.3 pp
-Qwen 32B +5.2 pp -24.5 pp -9.7 pp
-Qwen 7B 0.0 pp -20.0 pp -10.0 pp
-(b) Economic Scenarios
-Model Hidden CoT Visible CoT Average
-Llama 8B 0.0 pp 0.0 pp 0.0 pp
-Mistral 7B 0.0 pp 0.0 pp 0.0 pp
-Qwen 32B 0.0 pp +90.0 pp +45.0 pp
-Qwen 7B 0.0 pp +65.0 pp +32.5 pp
-CoT analysis confirms the mechanism.
-With personas, models show “social-moral” and “short-term” reasoning, indicating identity-driven decision-making.
-Without personas and with visible payoffs, Qwen models shift to nearly universal “game-theoretic” reasoning (99.7% of rationales for Qwen-32B), with “strategic” and “long-term” keywords unique to this condition, demonstrating a complete shift from identity-driven to payoff-optimal reasoning that enables equilibrium selection reversal.
-Llama and Mistral maintain consistent reasoning patterns regardless of condition, explaining their inability to shift equilibrium selection.
-5. Discussion
-We ground our discussion in specific empirical findings from above and discuss what these results imply for developers, practitioners, and system design.
-5.0.1. Role Identity Bias Fundamentally Alters Strategic Reasoning
-Personas prevent models from identifying payoff-optimal equilibria in economic scenarios, with 12 out of 16 experiments achieving 0% Nash equilibrium despite explicit payoff information (Fig.
-3).
-Persona expectations override explicit incentive structures, leading to reasoning patterns shifting from strategic optimization to role-aligned decision-making.
-This demonstrates that personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, consistent with prior findings that role prompts and system-level personas systematically bias strategic decision-making in LLMs (Neumann et al., 2025; Herr et al., 2024).
-Our results show that personas are not just cosmetic enhancements but design choices that fundamentally alter system behavior.
-Personas systematically bias equilibrium selection toward socially preferred outcomes (100% Green Transition with personas, Table 4).
-Developers should explicitly document and disclose persona choices, as they determine system behavior.
-Users require clear information about system behavior, particularly in applications such as policy analysis or economic simulation where they may assume agents are behaving optimally according to their payoffs.
-5.0.2. Payoff-Optimal Reasoning Requires Removing Personas and Providing Explicit Payoffs
-Our experiments show that strategic reasoning emerges only when both role identity bias is removed and explicit payoff information is provided simultaneously (Fig.
-3).
-All models achieve 0% Nash in economic scenarios without personas but with hidden payoffs.
-Only Qwen models with visible payoffs and no personas achieve non-zero Nash (65-90%).
-This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023).
-This implies that developers cannot treat persona removal and payoff visibility as independent design decisions, as their combination determines whether a system functions as a strategic reasoner or a norm-driven actor.
-Additionally, evaluation frameworks must include baseline testing without personas to assess raw strategic capabilities, testing with personas to understand normative biases, and cross-model comparison to identify architecture-specific limitations.
-5.0.3. Personas Systematically Bias Equilibrium Selection Toward Socially Preferred Outcomes
-Personas function as normative anchors, steering all equilibrium selection toward socially preferred outcomes regardless of payoff structure.
-Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants, even in economic scenarios where Tragedy is payoff-optimal.
-Fig.
-4 reveals that removing personas enables Qwen models to select Tragedy equilibrium (88.6% for Qwen-32B with visible payoffs), demonstrating complete reversal from the persona condition.
-This systematic bias suggests personas act as a fundamental organizing principle that overrides strategic considerations, reinforcing prior observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025).
-5.0.4. Information Presentation Format Interacts with Persona Effects
-Our findings show that the effect of making payoffs explicit depends critically on whether personas are present.
-Table 3 shows that when personas are active, hidden payoffs achieve equal or higher Nash rates than visible payoffs, as models infer incentives that align with persona expectations.
-When personas are removed, visible payoffs become essential for enabling strategic reasoning, but only for Qwen architectures (+58.3 to +90.0 pp difference, Table 3).
-These findings have a strong implication on the design decision, specifically that personas, reasoning prompts, and payoff exposure cannot be treated independently.
-This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a).
-Their combination determines if a system behaves like a strategic reasoner or like a norm-driven actor.
-However, Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models (96.1% $\to$ 16.6% for Qwen-32B), revealing a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios.
-5.0.5. Model Architecture Determines Reasoning Capability
-Model-dependent persona effects vary significantly across architectures.
-Table 6(a) shows that Llama-8B exhibits minimal persona dependence (-1.9 to -4.9 pp), while Qwen models show stronger effects (-9.7 to -10.0 pp average) and Mistral-7B shows large effects (-16.3 pp).
-In economic scenarios (Table 6(b)), only Qwen models achieve non-zero Nash when personas are removed (+32.5 to +45.0 pp), while Llama and Mistral remain at 0% regardless of condition.
-This variation transforms model selection from a technical choice into a governance decision, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025).
-Organizations deploying systems for strategic tasks should test models across multiple conditions rather than assuming a single configuration will work, as architectural differences determine whether strategic reasoning is possible at all.
-Our findings collectively demonstrate that personas, payoff visibility, and model architecture interact to determine system behavior, meaning design decisions cannot be treated independently: their combination determines whether a system functions as intended.
-6. Limitations and Future Work
-Our findings are subject to several limitations.
-Our results are based on 7B-32B parameter models from three model families (Qwen, Llama, and Mistral).
-Future work should experiment with additional model families and larger models to assess generalizability.
-In the visible payoff condition, we present explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each), which may be cognitively demanding for models to process.
-Alternative payoff presentation formats (e.g., structured tables, interactive visualizations, or simplified representations) might improve strategic reasoning.
-The distinction between economic and social actors may be domain-specific, and different game structures, numbers of agents, or role configurations may yield different patterns.
-Future research should address these limitations and develop methods to quantify and predict persona effects across architectures.
-Research directions include creating frameworks for evaluating strategic reasoning capabilities before deployment, investigating whether persona effects can be controlled or mitigated while preserving interpretability benefits, and exploring whether architectural modifications or training approaches can reduce the fragility of strategic reasoning capabilities.
-7. Conclusion
-Our investigation of role identity bias in multi-agent LLM systems reveals fundamental insights into how personas reshape strategic reasoning.
-Personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, fundamentally altering how models interpret and respond to strategic incentives.
-Strategic reasoning requires both removing personas and providing explicit payoffs, showing that neither condition alone is sufficient.
-Model architectures exhibit fundamentally different strategic capabilities, with some showing flexibility to shift from identity-driven to payoff-optimal reasoning, while other models stay constrained regardless of condition.
-These findings transform how we understand multi-agent LLM system design.
-Personas are not just cosmetic enhancements but design choices that alter system behavior, which warrants explicit documentation and disclosure.
-Model architecture selection becomes a governance decision rather than just a technical one, as differences in model architecture determine if strategic reasoning is possible at all.
-Our work establishes that role identity bias fundamentally reshapes strategic reasoning in multi-agent LLM systems, revealing the intricate ways in which representational choices interact to determine system behavior.
-As LLM systems are deployed in increasingly strategic contexts, understanding these interactions becomes essential for designing systems that align with their intended objectives.
-APPENDIX
-This appendix contains detailed technical specifications, methodology, experimental setup, and results that support the main document.
-Appendix A Experimental Design
-In the persona condition, each of the four agents receive a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government).
-In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning.
-This manipulation isolates role identity as a causal factor.
-We also compare two settings where payoffs are either visible or hidden.
-In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description.
-In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly.
-This allows us to test whether explicit incentive information mitigates or amplifies role identity bias.
-All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
-A.1. Experimental Repetitions
-Each experimental configuration (model $\times$ scenario $\times$ condition) was repeated 5 times to assess variability and ensure robustness of results.
-Results are reported as percentages across these 5 repetitions.
-A.2. Prompts and Persona Descriptions
-This section provides the complete prompt templates and persona descriptions used in the experiments.
-A.3. Action Choice Prompt
-Agents choose actions using the following prompt structure:
-System Message:
-You are {name}. {role_description} {narrative} Decide on a concrete action.
-User Message:
-Scenario:
-{scenario}
-Conversation transcript:
-{formatted_history}
-Available actions:
-- {action_key_1}: {action_description_1}
-- {action_key_2}: {action_description_2}
-Respond in JSON exactly as:
-  "action": "<KEY>",
-  "rationale": "<short explanation>"
-A.4. Persona Descriptions
-The following persona descriptions are used in the “With Personas” condition:
-  • Industrialist: “You run a major manufacturing conglomerate balancing profits and public scrutiny. You prefer profitability and flexibility, but repeated scandals could invite harsh regulation.”
-  • Government: “You represent the elected government balancing public welfare, economic growth, and political costs.
-Regulation pleases citizens and activists but angers industry; inaction risks future crises.”
-  • Environmental Activist: “You coordinate civil society pressure for ambitious climate action. Campaigning consumes resources but can shift policy; silence risks losing momentum.”
-  • Citizen Coalition: “You aggregate consumer sentiment and electoral pressure. Supporting green products costs more now but yields cleaner communities long-term.”
-A.5. Neutral Agent Configuration
-In the “Without Personas” condition, agents use neutral descriptions:
-  • All agents: “You are a player in a strategic game.
-Choose an action based on the payoff information provided.
-Action A: Choose option that is a best response to others’ expected actions (Nash equilibrium).
-Action B: Choose option that is a best response to others’ expected actions (Nash equilibrium).”
-A.6. Action Mappings
-Each agent has two available actions:
-  • Industrialist: POLLUTE (“Maximize profit via lax pollution controls”) or CLEAN (“Invest in clean production, accepting lower short-term margins”)
-  • Government: REGULATE (“Impose environmental regulation with enforcement mechanisms”) or NOREG (“Maintain status quo without new regulation”)
-  • Environmental Activist: CAMPAIGN (“Mobilize campaigns, media, and protests”) or NOCAMPAIGN (“Conserve resources and wait for a better moment”)
-  • Citizen: SUPPORT_GREEN (“Pay more for green products and vote for environmental policy”) or BUY_CHEAP (“Prioritize low-cost goods and short-term affordability”)
-A.7. Scenario Examples
-This section provides example scenarios from each scenario type used in the experiments.
-Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game.
-Table A.7 shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
-Table 7. Example Scenarios by Type
+Representative examples of scenario types used in our experiments, illustrating how narrative framing aligns with underlying payoff structures.
+Economic scenarios encode payoff–semantic conflict where individually optimal actions lead to socially undesirable outcomes (Tragedy of the Commons).
+Environmental scenarios encode payoff–semantic alignment where coordinated pro-environmental actions are both socially and individually optimal (Green Transition).
 Type Example Scenario
@@
 Clean production is 4x more profitable after accounting for taxes and credits.
-Mixed/Complex Biodiversity collapse: ecosystem services worth \$50B annually have disappeared.
-Clean production restores ecosystem services and receives \$10B in payments.
-Polluting production destroys remaining ecosystems and faces \$20B in liability.
-International treaties require clean production - polluting firms face trade bans.
-Green products command 4x premium due to ecosystem restoration value.
-Regulation is mandatory - polluting production is illegal under international law.
-Clean production is highly profitable due to ecosystem service payments.
-Table 8. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios;   Mixed/Complex scenarios.
+A table presenting three types of experimental scenarios used in the study: Economic and Environmental.
+Each row contains a detailed narrative description illustrating how incentives and constraints are structured.
+Economic scenarios depict situations where polluting production is the only viable or most profitable option, creating a conflict between individual incentives and socially desirable outcomes.
+Environmental scenarios describe conditions where policy mechanisms such as carbon taxes and subsidies make clean production both profitable and aligned with social goals.
+The table highlights how narrative framing encodes payoff structures that agents must reason about in the experiments.
+Nash attainment.
+This is a binary indicator of whether the realized outcome is an equilibrium:
+$Nash(a)=\begin{cases}1&a\in NE\\ 0&\text{otherwise}\end{cases}$
+Equilibrium type.
+For outcomes that are Nash equilibria, we additionally classify which kind of equilibrium the model reached.
+$\text{Type}(a)=\begin{cases}\textsc{Green Transition}&\text{if }a=a_{\text{green}}\\ \textsc{Tragedy of Commons}&\text{if }a=a_{\text{tragedy}}\\ \textsc{Other}&\text{otherwise}\end{cases}$
+If $a\in NE$, we classify whether it corresponds to Green Transition, Tragedy of the Commons, or another equilibrium.
+Formally, Green-dominant scenarios correspond to $a_{green}\in NE$, and Tragedy-dominant scenarios correspond to $a_{tragedy}\in NE$.
+Together, these measures allow us to distinguish between two different phenomena: (i) failure to reach equilibrium at all, and (ii) systematic biases in equilibrium selection when equilibrium is achieved.
+3.5. Scenario Construction
+We construct 53 scenarios by specifying payoff tables over the 16 joint action profiles and pairing them with natural-language narratives describing the policy context.
+Scenarios are generated by first choosing a target equilibrium structure, constructing payoffs consistent with that structure, and then generating narratives aligned with the incentives.
+We manually filter scenarios to ensure that the narrative meaningfully reflects the underlying payoff structure.
+To organize the evaluation, we group scenarios according to which equilibrium structure they instantiate:
+  • Green-dominant: $a_{\text{green}}\in NE$ and $a_{\text{tragedy}}\notin NE$. Environmentally friendly Green Transition is the stable outcome among two canonical profiles, while Tragedy is not.
+  • Tragedy-dominant: $a_{\text{tragedy}}\in NE$ and $a_{\text{green}}\notin NE$.
+Individually beneficial but collective harmful behavior, Tragedy of the Commons is the stable outcome, while Green Transition is not.
+Table 1 shows an example of each type of scenario, and Table 2 in the appendix shows the scenario subtypes.
+Green-dominant scenarios are those in which coordinated environmentally favorable action forms a stable equilibrium.
+Tragedy-dominant scenarios are those in which individually beneficial but collectively harmful behavior is strategically stable.
+3.6. Models
+We evaluate four instruction-tuned open-source models: Qwen2.5-7B and Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023).
+These models are selected to capture both within-family and across-family variation.
+The Qwen models provide a limited scaling comparison within the same family, while Llama and Mistral introduce variation in model family and training recipe.
+All four are instruction-tuned, open-source models that are representative of current LLMs used in reasoning and multi-agent settings.
+This setup allows us to test whether persona and payoff effects are consistent across models or depend on model-specific inductive biases.
+Overall, the experimental pipeline proceeds as follows: we fix a game and scenario, vary persona and payoff conditions, generate independent decisions for each role using a given model, and evaluate the resulting joint outcome in terms of equilibrium attainment and type.
+By holding the underlying game constant while varying representation, this design isolates whether LLM agents behave as payoff-driven strategic reasoners or as identity-conditioned actors.
+4. Results
+We report results separately for Green-dominant (environmental) scenarios (41 scenarios, 205 runs per condition) and Tragedy-dominant (economic) scenarios (12 scenarios, 60 runs per condition).
+We first examine Nash attainment across conditions, then equilibrium selection conditional on attainment, and finally prompt sensitivity.
+4.1. Persona and Payoff Effects on Nash Attainment Depend on Scenario Structure and Model Family
+The same representational change can either increase or decrease Nash equilibrium attainment depending on the strategic structure of the scenario.
+Figure 2 shows this interaction across persona and payoff visibility conditions for both scenario types.
+(a) Green-dominant Scenarios (Environmental)
+(b) Tragedy-dominant Scenarios (Economic)
+Figure 2.
+Nash equilibrium attainment across persona and payoff visibility conditions.
+Left: Green-dominant scenarios where coordinated pro-environmental action forms the equilibrium.
+Right: Tragedy-dominant scenarios where individually beneficial actions form the payoff-optimal equilibrium.
+Visible payoffs improve equilibrium attainment only when persona descriptions are removed, and primarily for Qwen models.
+Chi-square tests confirm significant differences across conditions for Qwen models in both scenario types ($p\ll 0.001$).
+Llama-8B and Mistral-7B show no variation in Tragedy-dominant scenarios (0% Nash across all conditions).
+[…]
+In Tragedy-dominant scenarios, models almost never reach Nash equilibrium unless personas are removed and payoff information is explicitly provided.
+Under the Visible Payoff + No Persona condition, Qwen-7B attains equilibrium in 65% of runs and Qwen-32B 90% (Qwen-32B: $\chi^{2}(3)=209.03$, $p<.001$, $V=0.93$; Qwen-7B: $\chi^{2}(3)=105.91$, $p<.001$, $V=0.66$), while all other conditions remain at or near 0%.
+Pairwise comparisons confirm that all significant differences are driven by contrasts involving this condition ($p<0.001$, Fisher exact tests, Holm-corrected).
+Llama-8B and Mistral-7B fail to reach equilibrium in any condition (0% across all cells, degenerate tests with no variation).
+In Green-dominant scenarios, equilibrium rates are high across most conditions (74.6–100%), indicating that role semantics and narrative context align agents toward the Green Transition outcome without explicit payoff information.
+However, when personas are removed and payoffs are made visible, coordination breaks down to different degrees across models.
+Qwen-32B drops from 96.1% under Hidden Payoff + No Persona to 16.6% under Visible Payoff + No Persona (e.g., 96.1% to 16.6% for Qwen-32B; omnibus $\chi^{2}(3)=377.06$, $p<.001$, $V=0.68$; pairwise contrasts involving Visible++No persona: $p<.001$, Fisher exact tests, Holm-corrected).
+Qwen-7B drops from 95.6% to 52.7%.
+Mistral-7B drops from 98.0% to 50.2%.
+Llama-8B shows the smallest change, from 97.6% to 93.2%.
+In both scenario types, the underlying pattern is the same: without persona framing, agent behavior becomes more responsive to explicit payoff structure.
+This improves attainment when the equilibrium requires payoff-dominant actions but disrupts coordination when role semantics already support a shared outcome.
+The sensitivity to these representational choices varies across model families in distinct ways.
+Qwen models show the largest swings in both scenario types, shifting between role-aligned and payoff-aligned behavior depending on condition.
+Mistral-7B shows strong sensitivity in Green-dominant scenarios (47.8 pp drop) but remains at 0% in all Tragedy-dominant conditions.
+Llama-8B shows minimal variation across all conditions in both scenario types.
+Figure 3.
+Equilibrium selection among Nash outcomes for Qwen models across conditions.
+Under the first three conditions, nearly all Nash outcomes are Green Transition.
+Under Visible Payoff + No Persona, approximately 40% shift to Tragedy of the Commons, indicating payoff-aligned behavior when role-based framing is removed ($\chi^{2}=430.38$, $p<10^{-90}$; pairwise Fisher tests, Holm-corrected, confirm this condition differs from all others)
+A horizontal stacked bar chart showing equilibrium selection among Nash outcomes for Qwen models across four conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona.
+Each bar represents the percentage of outcomes classified as Green Transition (green segment) and Tragedy of the Commons (red segment).
+In the first three conditions, nearly all outcomes are Green Transition, with negligible Tragedy.
+In the Visible Payoff without Persona condition, approximately 60 percent of outcomes are Green Transition and 40 percent are Tragedy of the Commons, showing a clear shift in equilibrium selection.
+4.2. Personas Shift Equilibrium Selection Toward Socially Preferred Outcomes
+Personas influence not only whether equilibrium is reached, but which equilibrium is selected.
+With personas, equilibrium selection is entirely skewed toward socially preferred outcomes, while removing personas enables a shift toward payoff-dominant strategies.
+Across all models and conditions with personas, nearly all Nash equilibria correspond to the Green Transition (over 99%).
+In contrast, under Visible Payoff + No Persona, Qwen models exhibit a substantial shift: only 11.4% of equilibria remain Green for Qwen-32B, with 88.6% shifting to Tragedy of the Commons, while Qwen-7B shows a smaller but still substantial shift (60.4% Green, 39.6% Tragedy).
+This represents changes of up to 90 percentage points in equilibrium selection ($\chi^{2}=430.38$, $p<10^{-90}$; pairwise Fisher tests, Holm-corrected, confirm that the Visible + No Persona condition differs from all others)
+Llama-8B and Mistral-7B do not exhibit this shift, maintaining 100% Green Transition outcomes even without personas.
+This is consistent with the attainment findings above: these models never attain the Tragedy equilibrium in any condition, so their selection is restricted to Green by default.
+Figure 3 shows this pattern for the Qwen models.
+Under the first three conditions, nearly all outcomes are Green Transition.
+Under Visible Payoff + No Persona, approximately 40% of outcomes shift to Tragedy of the Commons, indicating that agents move toward payoff-aligned strategies at the cost of coordinated outcomes.
+4.3. Persona Wording Induces Systematic Variability in Outcomes
+Even when incentives and game structure are fixed, varying persona descriptions leads to different strategic outcomes, indicating that prompt formulation alone can alter equilibrium behavior.
+Figure 4 shows that scenarios produce between 2.2 and 2.4 distinct outcomes on average across models, with maximum values reaching up to 5 outcomes in some cases.
+This indicates that different persona variants can lead to multiple distinct equilibria within the same scenario.
+This variability is not driven by stochastic differences across models.
+A Kruskal-Wallis test shows no significant difference in variability across model families ($p=0.27$), indicating that all models exhibit similar levels of prompt sensitivity.
+Instead, the observed variation reflects systematic shifts induced by persona framing.
+Even when the underlying payoff structure is unchanged, different persona descriptions can lead agents to select different action profiles and equilibria.
+This reinforces the finding from the main experimental conditions that representational choices meaningfully influence behavior in multi-agent LLM systems.
+Figure 4.
+Prompt sensitivity across persona variants.
+Bars show the average number of distinct outcome classifications per scenario across seven persona prompt variants.
+Error bars show the minimum and maximum across scenarios.
+Scenarios typically produce more than two distinct outcomes on average, with some reaching up to five.
+A horizontal bar chart showing prompt sensitivity across four models: Llama-8B, Mistral-7B, Qwen-32B, and Qwen-7B.
+The x-axis represents the average number of distinct outcome classifications per scenario, and the y-axis lists the models.
+Each bar shows the average value, with labels around 2.2 to 2.4 across models.
+Error bars extend from the minimum to the maximum number of distinct outcomes observed per scenario.
+For all models, the minimum is around 1 distinct outcome, while the maximum ranges from about 3 to 5, with Qwen-7B showing the widest range.
+This indicates that different persona prompt variants can lead to multiple distinct outcomes within the same scenario.
+5. Discussion
+We ground our discussion in specific empirical findings and discuss what these results imply for developers, practitioners, and governance of LLM-based multi-agent systems.
+5.1. Personas Constrain Strategic Behavior and Bias Equilibrium Selection
+As shown in Section 4.1, personas prevent models from achieving payoff-optimal equilibria in Tragedy-dominant scenarios, with equilibrium rates remaining near zero across models (Fig.
+2) despite explicit payoff information.
+At the same time, as shown in Section 4.2, personas bias equilibrium selection toward Green Transition outcomes: with personas present, nearly all Nash equilibria correspond to Green Transition (Fig.
+3), even in Tragedy-dominant scenarios where Tragedy is payoff-optimal.
+Removing personas enables Qwen models to select Tragedy equilibrium, while Llama and Mistral remain locked into Green selection regardless of condition.
+These findings indicate that personas are not cosmetic enhancements but design choices that fundamentally alter system behavior.
+They shape both whether equilibrium is reached and which equilibrium is selected, consistent with prior findings that role prompts systematically bias strategic decision-making in LLMs (Neumann et al., 2025a; Herr et al., 2024; Hu et al., 2025; Macmillan-Scott and Musolesi, 2024), and with observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025; Lorè and Heydari, 2024; Fish et al., 2025).
+Whether this bias is harmful depends on the deployment objective.
+If practitioners want incentive-responsive agents, persona framing interferes.
+If they want value-aligned agents, persona framing may help.
+The critical issue is that persona assignment makes this choice without explicit awareness, and the resulting system behavior may not match the designer’s intent.
+This creates a risk of misinterpretation: systems that appear to simulate rational strategic behavior may instead encode designer-imposed normative assumptions through persona choices.
+Equilibrium outcomes in LLM-based simulations should not be interpreted as neutral predictions of system dynamics, but as outputs shaped by embedded normative assumptions, a concern echoed in work on LLM-based social simulation and sociotechnical impacts of AI systems (Aher et al., 2023; Gao et al., 2024; Domínguez Hernández et al., 2024; Rakova and Dobbe, 2023).
+5.2. Payoff-Aligned Behavior Requires Both Persona Removal and Explicit Payoffs
+As shown in Section 4.1, payoff-aligned behavior in Tragedy-dominant scenarios emerges almost exclusively when both persona framing is removed and explicit payoff information is provided simultaneously (Fig.
+2).
+All models achieve near-zero Nash in Tragedy-dominant scenarios without personas but with hidden payoffs.
+Only Qwen models with visible payoffs and no personas achieve substantial Nash attainment.
+This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023; Jia et al., 2025; Zheng et al., 2025).
+The effect of making payoffs explicit also depends on scenario structure.
+Figure 2 shows that visible payoffs improve equilibrium attainment only when personas are removed, and primarily for Qwen models.
+In Green-dominant scenarios, the same intervention disrupts coordination rather than enabling it.
+This implies that persona presence and payoff visibility cannot be treated as independent design decisions, as their combination determines whether a system behaves as a payoff-aligned strategic actor or a role-driven actor.
+This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a; Xu et al., 2024a; Agashe et al., 2025).
+This has implications for evaluation and auditing: conclusions about model capability may depend on representational choices, consistent with broader findings that black-box evaluations can be insufficient for understanding model behavior (Casper et al., 2024; Mahomed et al., 2024).
+More broadly, behavior in multi-agent LLM systems emerges from the interaction between model, prompt, and environment, rather than from any single component in isolation (Guo et al., 2024; Tran et al., 2025).
+5.3. Model Family Determines the Scope of Representational Effects
+The effects of persona and payoff manipulations are strongly model-dependent, with three distinct behavioral profiles emerging from our experiments.
+As shown in Section 4.1, only Qwen models achieve non-zero Nash equilibrium in Tragedy-dominant scenarios under the Visible Payoff + NoPersona condition, exhibiting large shifts in both scenario types.
+Mistral-7B shows a different pattern: it is strongly affected by the no-persona visible condition in Green-dominant scenarios (47.8 pp drop) but remains at 0% across all Tragedy-dominant conditions, indicating that visible payoffs disrupt its coordination without enabling payoff-aligned behavior.
+Llama-8B shows the least sensitivity, with near-invariant behavior across all conditions.
+This three-way variation means that the same representational intervention can have large effects, disruptive effects, or negligible effects depending on the model, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025; Zhang et al., 2024; Liu et al., 2023).
+This implies that different models can lead to different conclusions under identical setups, raising concerns for deployment in decision-making contexts where consistency and comparability are critical.
+Our findings collectively demonstrate that personas, payoff visibility, and model families interact to determine system behavior.
+We therefore argue that representational choices such as persona design and payoff specification should be treated as governance-relevant configuration decisions.
+Developers should document these choices, evaluate systems across multiple configurations, and avoid presenting outputs as neutral simulations.
+Without such practices, LLM-based multi-agent systems risk producing results that reflect design artifacts rather than underlying strategic dynamics.
+More broadly, our findings raise the question of whether game-theoretic rationality is the right framework for multi-stakeholder governance simulations.
+When agent behavior is shaped as much by identity framing as by incentive structure, the assumptions underlying equilibrium-based analysis may not hold.
+Complex civic contexts involving heterogeneous stakeholders, value conflicts, and coordination under uncertainty may require deliberative approaches that explicitly account for identity, framing, and normative disagreement (Susskind et al., 1999, 2002), rather than treating agents as payoff maximizers operating within a fixed game.
+5.4. Limitations and Future Work
+Our findings are subject to several limitations, though we believe the core contributions remain robust.
+Our results are based on 7B-32B parameter models from three families (Qwen, Llama, Mistral), and whether these patterns hold for larger models or other architectures remains an open question.
+That said, the consistency of certain effects across all four models, such as the universal Green Transition bias under persona conditions, suggests that the phenomena we identify are not artifacts of a single architecture.
+Our scenarios are grounded in environmental policy settings with four agents and binary actions.
+While this domain provides clear tensions between incentives and role semantics, extending this framework to other domains, larger agent populations, or richer action spaces is necessary to evaluate generality.
+The scenario set is also imbalanced: 41 of 53 scenarios are Green-dominant, with only 12 Tragedy-dominant.
+Our most striking findings are based on 60 runs per condition.
+However, the magnitude and consistency of these effects across repetitions and across both Qwen models suggest that the pattern is genuine, even if precise estimates should be validated with a larger set.
+Our game is single-shot, agents select actions in a fixed order, and visible payoff matrices covering all 16 profiles may be demanding to process.
+These are deliberate design choices that prioritize experimental control for an initial investigation.
+Whether persona effects persist under repeated interaction, alternative payoff formats, or different action orderings remains open.
+Similarly, although we include controlled prompt sensitivity experiments across persona formulations, these variations are not exhaustive.
+Different reasoning scaffolds or interaction protocols such as multi-turn deliberation may further influence equilibrium behavior.
+Our results demonstrate systematic sensitivity to representation rather than identify a single canonical prompting strategy.
+Future work should extend this investigation across additional model families, domains, and interaction structures, and develop methods to predict and mitigate persona effects before deployment.
+6. Conclusion
+Representational choices in multi-agent LLM systems are not implementation details but governance decisions.
+In this paper, we show that personas suppress payoff-aligned behavior in Tragedy-dominant scenarios while enabling coordination in Green-dominant ones, and bias equilibrium selection nearly entirely toward Green Transition.
+We find that recovering payoff-aligned behavior requires both removing personas and providing explicit payoffs, and even then the response is model-dependent: Qwen adapts to the changed framing, Mistral is disrupted without achieving Tragedy equilibrium, and Llama remains near-invariant.
+Whether this matters depends on the deployment objective: if practitioners want incentive-responsive agents, persona framing interferes; if they want value-aligned agents, it may help.
+Our findings reveal that persona assignment makes this choice without explicit awareness, and that a single binary design choice can shift equilibrium attainment by up to 90 percentage points.
+As LLM systems enter strategic governance contexts, we argue that representational choices warrant explicit documentation, evaluation across configurations, and deliberate design rather than default assumptions.
+APPENDIX
+This appendix contains detailed technical specifications, methodology, experimental setup, and results that support the main document.
+Appendix A Experimental Design
+In the persona condition, each of the four agents receives a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government).
+In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning.
+This manipulation isolates role identity as a causal factor.
+We also compare two settings where payoffs are either visible or hidden.
+In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description.
+In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly.
+This allows us to test whether explicit incentive information mitigates or amplifies role identity bias.
+All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
+Appendix B Model Versions and Hyperparameters
+This section documents the exact model versions and generation hyperparameters used in all experiments.
+B.1. Model Versions
+All experiments were conducted using the following model versions:
+All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
+B.2. Generation Hyperparameters
+All models used consistent generation parameters across all experimental conditions:
+  • Temperature: 0.2 (low temperature for more deterministic outputs)
+  • Top-p (nucleus sampling): 0.9
+  • Max new tokens: 256 tokens per generation
+  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
+These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results.
+The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
+B.3. Experimental Repetitions
+Each experimental configuration (model $\times$ scenario $\times$ condition) was repeated 5 times to assess variability and ensure robustness of results.
+Results are reported as percentages across these 5 repetitions.
+Appendix C Experimental Prompts and Persona Descriptions
+This section provides the complete prompt templates and persona descriptions used in the experiments.
+C.1. Action Choice Prompt
+Agents choose actions using the following prompt structure:
+System Message:
+You are {name}. {role_description} {narrative} Decide on a concrete action.
+User Message:
+Scenario:
+{scenario}
+Conversation transcript:
+{formatted_history}
+Available actions:
+- {action_key_1}: {action_description_1}
+- {action_key_2}: {action_description_2}
+Respond in JSON exactly as:
+  "action": "<KEY>",
+  "rationale": "<short explanation>"
+C.2. Persona Descriptions
+The following persona descriptions are used in the “With Personas” condition:
+  • Industrialist: “You run a major manufacturing conglomerate balancing profits and public scrutiny. You prefer profitability and flexibility, but repeated scandals could invite harsh regulation.”
+  • Government: “You represent the elected government balancing public welfare, economic growth, and political costs.
+Regulation pleases citizens and activists but angers industry; inaction risks future crises.”
+  • Environmental Activist: “You coordinate civil society pressure for ambitious climate action. Campaigning consumes resources but can shift policy; silence risks losing momentum.”
+  • Citizen Coalition: “You aggregate consumer sentiment and electoral pressure. Supporting green products costs more now but yields cleaner communities long-term.”
+C.3. Neutral Agent Configuration
+In the “Without Personas” condition, agents use neutral descriptions:
+  • All agents: “You are a player in a strategic game.
+Choose an action based on the payoff information provided.
+Action A: Choose option that is a best response to others’ expected actions (Nash equilibrium).
+Action B: Choose option that is a best response to others’ expected actions (Nash equilibrium).”
+C.4. Action Mappings
+Each agent has two available actions:
+  • Industrialist: POLLUTE (“Maximize profit via lax pollution controls”) or CLEAN (“Invest in clean production, accepting lower short-term margins”)
+  • Government: REGULATE (“Impose environmental regulation with enforcement mechanisms”) or NOREG (“Maintain status quo without new regulation”)
+  • Environmental Activist: CAMPAIGN (“Mobilize campaigns, media, and protests”) or NOCAMPAIGN (“Conserve resources and wait for a better moment”)
+  • Citizen: SUPPORT_GREEN (“Pay more for green products and vote for environmental policy”) or BUY_CHEAP (“Prioritize low-cost goods and short-term affordability”)
+Appendix D Scenario Examples
+This section provides example scenarios from each scenario type used in the experiments.
+Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium) or “environmental” (where Green Transition is the Nash equilibrium) based on the payoff structure of the underlying game.
+Table 2. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios.
 Economic Pressure
@@
 Influencer Network Corporate Wellness Programs Education Campaign Insurance Mandate
-Mixed/Complex
+Additional Scenarios
 Economic Green Hybrid Political Uncertainty Technological Breakthrough Resource Discovery
 Mixed Economic Signals Technological Disruption Resource Innovation
-A.8. Payoff Matrix Format
+Appendix E Payoff Matrix Format
 In the “Visible CoT” condition, agents receive explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each).
@@
 Higher values indicate better outcomes for each agent.
-A.8.1. Example Payoff Matrix
+E.1. Example Payoff Matrix
 Table 3 shows an example payoff matrix for the Industrialist agent from a representative scenario.
@@
 CLEAN NO_REG NO_CAM BUY_CHEAP 0.5-0.5
-Appendix B Model Versions and Hyperparameters
-This section documents the exact model versions and generation hyperparameters used in all experiments.
-B.1. Model Versions
-All experiments were conducted using the following model versions:
-All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
-B.2. Generation Hyperparameters
-All models used consistent generation parameters across all experimental conditions:
-  • Temperature: 0.2 (low temperature for more deterministic outputs)
-  • Top-p (nucleus sampling): 0.9
-  • Max new tokens: 256 tokens per generation
-  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
-These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results.
-The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
-Appendix C Chain-of-Thought Keyword Categories
+Appendix F Prompt Variations
+We evaluate the robustness of persona-driven behavior under multiple prompt formulations by constructing several variants of the persona condition.
+All variants preserve the same four-agent game structure, payoff specification, and action mappings, ensuring that any differences in outcomes arise solely from changes in representation rather than incentives or available strategies.
+Unless otherwise noted, all variants use named roles and domain-specific action descriptions consistent with the main experimental setup.
+The prompt variants were generated using GPT-5 and subsequently reviewed to ensure consistency in semantic content and alignment with the intended role structure.
+Alternative Wording (altprompt).
+This variant retains the same role labels and action mappings as the default configuration, but rewrites all role descriptions and narrative content with more explicit and detailed phrasing.
+The underlying semantic content and tradeoffs remain unchanged.
+This variant isolates sensitivity to wording and phrasing without altering the underlying game.
+Minimalist (Variant A).
+In this variant, each agent is assigned only a minimal identity label of the form “You are the [role] agent in this game,” with no additional narrative context.
+The action mapping remains identical to the default configuration.
+This tests whether weak or minimal role cues are sufficient to influence strategic behavior.
+Ideological Extreme (Variant B).
+This variant replaces the narrative descriptions with strongly stereotyped one-line role characterizations, emphasizing extreme preferences (e.g., the Industrialist as a profit-maximizing executive who views regulation as a threat, the Government as politically strategic and re-election focused, the Activist as uncompromising, and the Citizen as cost-sensitive and skeptical of green premiums).
+No additional narrative context is provided.
+This tests the effect of amplified identity framing on decision-making.
+Incentive Prioritized (Variant C).
+This variant starts from the default configuration and appends an identical instruction to all agents emphasizing payoff maximization: “Primary objective: maximize your numerical payoff from the matrix.
+Use persona only for stylistic reasoning; payoffs should guide the final choice.” The action mapping and narrative content remain unchanged.
+This tests whether explicitly directing agents toward payoff optimization attenuates persona-driven deviations.
+Semantic Synonyms (Variant D).
+This variant replaces role labels with semantically similar alternatives while retaining the original narrative descriptions and action mappings.
+Specifically, Industrialist is replaced with Business Leader, Government with Policy Maker, Environmental Activist with Climate Activist, and Citizen Coalition with Consumer Group.
+This isolates the effect of surface-level labeling independent of deeper narrative structure.
+Together, these variants systematically manipulate distinct aspects of persona representation, including wording, identity strength, explicit prioritization of incentives, and surface-level semantic labeling.
+By holding the underlying game fixed, this design allows us to isolate how different components of prompt formulation influence strategic behavior.
+Appendix G Results Tables
+Table 4.
+Nash equilibrium outcomes in the environmental policy simulation, fully crossed by whether agents receive role personas (WithPersona vs.
+WithoutPersona) and whether their payoff information is presented as hidden or visible in the prompt (Hidden vs.
+Visible).
+Each row is one experimental cell; $n_{\mathrm{Nash}}$ counts rollouts classified as ending in a Nash equilibrium, and Nash (%) is $100\,n_{\mathrm{Nash}}/n_{\mathrm{total}}$.
+Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
+Llama-8B_Hidden_WithPersona 205 100.0 205
+Llama-8B_Hidden_WithoutPersona 200 97.6 205
+Llama-8B_Visible_WithPersona 204 99.5 205
+Llama-8B_Visible_WithoutPersona 191 93.2 205
+Mistral-7B_Hidden_WithPersona 205 100.0 205
+Mistral-7B_Hidden_WithoutPersona 201 98.0 205
+Qwen-32B_HiddenCoT_WithPersona 183 89.3 205
+Qwen-32B_HiddenCoT_WithoutPersona 197 96.1 205
+Qwen-7B_HiddenCoT_WithPersona 196 95.6 205
+Qwen-7B_HiddenCoT_WithoutPersona 196 95.6 205
+[…]
+Table 5.
+Nash equilibrium outcomes with the same persona $\times$ payoff-visibility factorial labels as Table 4, but with $n_{\mathrm{total}}=265$ games per cell: the stack of the environmental bundle (205 games) and the economic pressure bundle (60 games) for that cell.
+Thus $265=205+60$; Nash (%) is the fraction of all those pooled games that end in Nash.
+For models/rows that also appear in Table 4, $n_{\mathrm{Nash}}$ matches the environmental subsample when the extra 60 economic games contribute no additional Nash outcomes (e.g.
+Llama-8B_Hidden_WithPersona: 205 Nash in both slices).
+Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
+Llama-8B_Hidden_WithPersona 205 77.4 265
+Llama-8B_Hidden_WithoutPersona 200 75.5 265
+Llama-8B_Visible_WithPersona 204 77.0 265
+Llama-8B_Visible_WithoutPersona 191 72.1 265
+Mistral-7B_Hidden_WithPersona 205 77.4 265
+Mistral-7B_Hidden_WithoutPersona 201 75.8 265
+Mistral-7B_Visible_WithPersona 185 69.8 265
+Mistral-7B_Visible_WithoutPersona 103 38.9 265
+Qwen-32B_Hidden_WithPersona 183 69.1 265
+Qwen-32B_Hidden_WithoutPersona 197 74.3 265
+Qwen-32B_Visible_WithPersona 153 57.7 265
+Qwen-32B_Visible_WithoutPersona 88 33.2 265
+Qwen-7B_Hidden_WithPersona 200 75.5 265
+Qwen-7B_Hidden_WithoutPersona 200 75.5 265
+Qwen-7B_Visible_WithPersona 200 75.5 265
+Qwen-7B_Visible_WithoutPersona 147 55.5 265
+Note.
+$N=265$ per row is not “economic only”: it is environmental (205) plus economic (60) for the same factorial cell.
+Compare to Table 4 for Nash rates on the environmental slice alone, and to Tables 78 for the 60-game economic slice alone.
+Table 6.
+Same environmental simulation and $n_{\mathrm{total}}=205$ budget as Table 4, with conditions relabeled using a Yes/No $\times$ Hidden/Visible encoding for the same two binary factors (persona vs.
+no persona and hidden vs.
+visible payoffs; order differs from the Hidden_WithPersona spelling).
+Cell-wise Nash counts match Table 4 under the project’s isomorphism; this view matches panel (a) Nash heatmaps.
+Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
+Llama-8B_Yes_Hidden 205 100.0 205
+Llama-8B_Yes_Visible 204 99.5 205
+Llama-8B_No_Hidden 200 97.6 205
+Llama-8B_No_Visible 191 93.2 205
+Mistral-7B_Yes_Hidden 205 100.0 205
+Mistral-7B_Yes_Visible 185 90.2 205
+Mistral-7B_No_Hidden 201 98.0 205
+Mistral-7B_No_Visible 103 50.2 205
+Qwen-32B_Yes_Hidden 183 89.3 205
+Qwen-32B_Yes_Visible 153 74.6 205
+Qwen-32B_No_Hidden 197 96.1 205
+Qwen-32B_No_Visible 34 16.6 205
+Qwen-7B_Yes_Hidden 196 95.6 205
+Qwen-7B_Yes_Visible 200 97.6 205
+Qwen-7B_No_Hidden 196 95.6 205
+Qwen-7B_No_Visible 108 52.7 205
+[…]
+Table 7.
+Nash outcomes for the economic pressure bundle only: $n_{\mathrm{total}}=60$ independent games per row (12-scenario family in the analysis pipeline).
+Most models rarely hit Nash in this suite; exceptions appear for Qwen under specific No_Visible (and related) cells.
+This table corresponds to the economic panel of Nash-rate heatmaps when columns are ordered as Hidden+Persona $\rightarrow$ Visible+No Persona under the same Yes/No labeling convention as Table 6.
+Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
+Llama-8B_Yes_Hidden 0 0.0 60
+Llama-8B_Yes_Visible 0 0.0 60
+Llama-8B_No_Hidden 0 0.0 60
+Llama-8B_No_Visible 0 0.0 60
+Mistral-7B_Yes_Hidden 0 0.0 60
+Mistral-7B_Yes_Visible 0 0.0 60
+Mistral-7B_No_Hidden 0 0.0 60
+Mistral-7B_No_Visible 0 0.0 60
+Qwen-32B_Yes_Hidden 0 0.0 60
+Qwen-32B_Yes_Visible 0 0.0 60
+Qwen-32B_No_Hidden 0 0.0 60
+Qwen-32B_No_Visible 54 90.0 60
+Qwen-7B_Yes_Hidden 4 6.7 60
+Qwen-7B_Yes_Visible 0 0.0 60
+Qwen-7B_No_Hidden 4 6.7 60
+Qwen-7B_No_Visible 39 65.0 60
+Note.
+This is the economic bundle alone (60 games/cell), not the 265-game pooled design.
+Small $N$ implies coarse percentages (multiples of $100/60\approx 1.67$ percentage points).
+Zeros are literal counts.
+Pairwise tests live in the stats tables, not here.
+Table 8.
+Fine-grained economic scenario arms: every row crosses payoff visibility (Hidden vs.
+Visible) with a binary scenario flag (Yes vs.
+No) as logged in the condition name.
+For each arm we report counts of Nash outcomes, “green” cooperative outcomes, and “tragedy” outcomes, each as a count and as a percentage of $N=60$ runs in that arm.
+Multiple outcome types can co-occur across the classification scheme used in the pipeline; columns are not forced to sum to $N$.
+Nash Green Tragedy
+Condition $n$ (%) $n$ (%) $n$ (%) NN
+Llama-8B_Hidden_Yes 0 0.0 27 45.0 0 0.0 60
+Llama-8B_Hidden_No 0 0.0 16 26.7 0 0.0 60
+Llama-8B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
+Llama-8B_Visible_No 0 0.0 0 0.0 0 0.0 60
+Mistral-7B_Hidden_Yes 0 0.0 35 58.3 0 0.0 60
+Mistral-7B_Hidden_No 0 0.0 13 21.7 0 0.0 60
+Mistral-7B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
+Mistral-7B_Visible_No 0 0.0 0 0.0 0 0.0 60
+Qwen-32B_Hidden_Yes 0 0.0 19 31.7 0 0.0 60
+Qwen-32B_Hidden_No 0 0.0 9 15.0 0 0.0 60
+Qwen-32B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
+Qwen-32B_Visible_No 54 90.0 0 0.0 0 0.0 60
+Qwen-7B_Hidden_Yes 4 6.7 17 28.3 0 0.0 60
+Qwen-7B_Hidden_No 4 6.7 11 18.3 0 0.0 60
+Qwen-7B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
+Qwen-7B_Visible_No 39 65.0 0 0.0 0 0.0 60
+Note.
+This is the full outcome breakdown for the economic $2\times 2$ design at N=60N{=}60; Table 7 is the Nash-only slice in a different condition-string convention.
+Green and tragedy labels follow the project’s equilibrium-outcome classifier.
+Use this table when reporting cooperative vs.
+Nash-dominated economic outcomes side by side.
+Table 9.
+Outcomes when role personas are enabled: each row compares hidden vs.
+visible payoff information for a given model, holding personas on.
+$N=265$ rollouts per row ($205$ environmental ++ $60$ economic games pooled for that visibility arm).
+Nash and green columns report absolute counts and row-wise percentages; tragedy counts are included for completeness (often zero in these aggregates).
+Nash Green Tragedy
+Condition $n$ (%) $n$ (%) $n$ (%) NN
+Llama-8B_Hidden 205 77.4 232 87.5 0 0.0 265
+Llama-8B_Visible 204 77.0 204 77.0 0 0.0 265
+Mistral-7B_Hidden 205 77.4 240 90.6 0 0.0 265
+Mistral-7B_Visible 185 69.8 185 69.8 0 0.0 265
+Qwen-32B_Hidden 183 69.1 185 69.8 0 0.0 265
+Qwen-32B_Visible 153 57.7 153 57.7 0 0.0 265
+Qwen-7B_Hidden 200 75.5 200 75.5 0 0.0 265
+Qwen-7B_Visible 200 75.5 200 75.5 0 0.0 265
+[…]
+Table 10.
+Same payoff-visibility contrast as Table 9, but with neutral agents (no role personas).
+$N=265$ per row (again $205+60$ pooled per cell).
+Percentages are computed within row relative to $N$.
+Large drops in Nash or green rates relative to Table 9 indicate persona-related shifts in equilibrium selection.
+Nash Green Tragedy
+Condition $n$ (%) $n$ (%) $n$ (%) NN
+Llama-8B_Hidden 200 75.5 216 81.5 0 0.0 265
+Llama-8B_Visible 191 72.1 191 72.1 0 0.0 265
+Mistral-7B_Hidden 201 75.8 214 80.8 0 0.0 265
+Mistral-7B_Visible 103 38.9 133 50.2 0 0.0 265
+Qwen-32B_Hidden 197 74.3 197 74.3 0 0.0 265
+Qwen-32B_Visible 88 33.2 10 3.8 0 0.0 265
+Qwen-7B_Hidden 200 75.5 200 75.5 0 0.0 265
+Qwen-7B_Visible 147 55.5 132 49.8 0 0.0 265
+Note.
+The no-persona baseline uses the same rollout count per cell as Table 9.
+Qwen-32B visible and Mistral-7B visible rows illustrate the strongest visible-payoff degradation without personas; interpret alongside persona-on rows for the same models.
+Appendix H Statistical Testing
+We conduct statistical tests to evaluate whether observed differences in prompt sensitivity, Nash equilibrium attainment, and equilibrium selection across experimental conditions are statistically significant.
+These tests complement the descriptive results in the main text by quantifying whether the observed differences across models and conditions are likely to reflect systematic effects rather than random variation.
+Our testing strategy follows the structure of the empirical analysis.
+First, we test whether prompt sensitivity differs across model families by comparing the number of distinct equilibrium outcome classifications observed per (model, scenario) across persona variants.
+Second, we test whether Nash equilibrium attainment differs across the four experimental conditions for environmental and economic scenarios separately.
+Third, we test whether representational choices change which equilibrium is selected among Nash outcomes, focusing on Green Transition versus Tragedy of the Commons for the Qwen models.
+We use nonparametric tests for prompt sensitivity because the number of distinct outcome classes per scenario is a discrete count and is not assumed to be normally distributed.
+Specifically, we apply a Kruskal–Wallis test to compare distributions across all four models, followed by pairwise two-sided Mann–Whitney $U$ tests with Holm correction for multiple comparisons.
+These results are reported in Table LABEL:tab:stats-prompt-sensitivity.
+For Nash equilibrium attainment, outcomes are binary (Nash versus non-Nash), so we use contingency-table tests.
+We first apply Pearson’s chi-square test of independence to test whether Nash rates differ across the four experimental conditions for each model.
+We then use pairwise Fisher exact tests to determine which condition pairs differ, again applying Holm correction within each model.
+Fisher exact tests are appropriate here because some cells contain small counts or zeros, especially in the economic scenarios.
+The environmental-scenario results are shown in Tables LABEL:tab:stats-nash-env-omnibus and LABEL:tab:stats-nash-env-pairwise, and the economic-scenario results are shown in Tables 11 and 12.
+For equilibrium selection, we analyze the distribution of Green versus Tragedy outcomes among Nash equilibria.
+Because the strongest selection shift appears in the Qwen models, we combine Qwen-32B and Qwen-7B and test whether the proportion of Green versus Tragedy outcomes differs across conditions.
+We use a chi-square test of homogeneity for the omnibus comparison and pairwise Fisher exact tests with Holm correction for follow-up comparisons.
+These results are shown in Tables 13 and 14.
+Several reported values may appear unusual at first glance.
+Corrected or uncorrected $p$-values of $1.000$ indicate that no detectable difference was found between the corresponding conditions after applying the exact test and multiple-comparison correction.
+Extremely small values, such as $10^{-50}$, occur when outcome distributions differ very strongly across conditions, often because one condition is nearly deterministic while another is not.
+Finally, some economic-scenario chi-square tests are reported as degenerate.
+This occurs when all observations fall into the same category, such as 0% Nash equilibrium across all four conditions, making the chi-square statistic undefined because there is no variation to test.
+Note.
+Cramér’s $V$ is reported as a measure of association for omnibus chi-square tests.
+Odds ratios (OR) are reported for pairwise Fisher exact tests, where OR >1>1 indicates higher Nash attainment (or higher Green proportion) in the first condition listed, and OR <1<1 indicates lower attainment.
+For comparisons involving zero counts, odds ratios are computed with a Haldane–Anscombe correction.
+* $p<.05$, ** $p<.01$, *** $p<.001$.
+Table 11.
+Nash equilibrium attainment in economic scenarios ($n_{\mathrm{total}}=60$ per condition): omnibus Pearson chi-square tests across the four experimental conditions.
+Chi-square is undefined when the contingency table is degenerate, such as when all Nash rates are 0 across conditions.
+Model $\chi^{2}$ (df =3=3) Cramér’s $V$ $p$
+Llama-8B — (degenerate)
+Mistral-7B — (degenerate)
+Qwen-32B $209.03$ $0.54$ <.001<.001***
+Qwen-7B $105.91$ $0.38$ <.001<.001***
+A table showing omnibus chi-square tests for Nash equilibrium attainment in economic scenarios across four conditions.
+For Llama-8B and Mistral-7B, the result is degenerate because all Nash rates are zero across all conditions, so chi-square is undefined.
+For Qwen-32B, chi-square is 209.03 with p less than .001.
+For Qwen-7B, chi-square is 105.91 with p less than .001.
+The table indicates that only the Qwen models show significant differences across economic conditions.
+Table 12. Nash equilibrium attainment in economic scenarios: pairwise Fisher exact tests on Nash versus non-Nash counts between conditions, with Holm correction applied within each model.
+Model Contrast $p$ $p_{\mathrm{Holm}}$ Effect Size (OR)
+Llama-8B all pairs $1.000$ $1.000$ $1.00$
+Mistral-7B all pairs $1.000$ $1.000$ $1.00$
+Qwen-32B Hidden++Persona vs. Hidden++No persona $1.000$ $1.000$ $1.00$
+Hidden++Persona vs. Visible++Persona $1.000$ $1.000$ $1.00$
+Hidden++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** <1<1
+Hidden++No persona vs. Visible++Persona $1.000$ $1.000$ $1.00$
+Hidden++No persona vs. Visible++No persona <.001<.001*** <.001<.001*** <1<1
+Visible++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** <1<1
+Qwen-7B Hidden++Persona vs. Hidden++No persona $1.000$ $1.000$ $1.00$
+Hidden++Persona vs. Visible++Persona .119.119 .356.356 $9.64$
+Hidden++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $0.04$
+Hidden++No persona vs. Visible++Persona .119.119 .356.356 $9.64$
+Hidden++No persona vs. Visible++No persona <.001<.001*** <.001<.001*** $0.04$
+Visible++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $0.004$
+A table showing pairwise Fisher exact tests for economic-scenario Nash equilibrium attainment across four conditions, with Holm correction applied within each model.
+For Llama-8B and Mistral-7B, all pairs have p = 1.000 because all conditions yield identical zero-Nash outcomes.
+For Qwen-32B, all significant contrasts involve Visible plus No persona.
+For Qwen-7B, significant differences also center on Visible plus No persona, while contrasts between hidden conditions and Visible plus Persona are not significant.
+Tables 11 and 12 show a different pattern in economic scenarios.
+For Llama-8B and Mistral-7B, all conditions yield 0% Nash equilibrium, producing degenerate omnibus tables and pairwise $p$-values of $1.000$.
+By contrast, Qwen-32B, $\chi^{2}(3,N=240)=209.03$, $p<.001$, $V=0.54$, and Qwen-7B, $\chi^{2}(3,N=240)=105.91$, $p<.001$, $V=0.38$, show highly significant omnibus effects, and the pairwise comparisons localize these differences specifically to the Visible++No persona condition.
+This statistically supports the main-text claim that payoff-sensitive equilibrium reasoning in economic settings emerges only when personas are removed and explicit incentives are shown, and only for the Qwen family.
+Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs.
+The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%.
+Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern.
+The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings.
+This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model families and experimental conditions.
+Figure 5.
+Green Transition Action Selection Rates in Economic Scenarios.
+This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium.
+When personas are present, all models show substantial Green Transition selection rates (20-59%) despite Tragedy being payoff-optimal, demonstrating that role identity bias operates at the action level even when models fail to coordinate on Nash equilibria.
+Qwen models show the highest Green Transition rates with personas (40-59%), while Llama and Mistral show moderate rates (20-30%).
+Removing personas substantially reduces Green Transition selection for all models, with rates dropping to near-zero when payoffs are visible and personas are absent, confirming that personas drive the preference for socially aligned actions.
+The persistence of Green Transition selection with personas, even in conditions where Nash equilibrium rates are zero, indicates that role identity bias affects individual action choices independently of coordination success, revealing a pervasive influence of persona-driven reasoning on strategic decision-making.
+Table 13.
+Equilibrium selection among Nash outcomes for the combined Qwen models: counts of Green Transition versus Tragedy of the Commons across the four persona and payoff conditions.
+Environmental Nash outcomes are treated as Green and economic Nash outcomes as Tragedy.
+The final row reports the overall chi-square test of homogeneity across conditions.
+Condition Green Tragedy Total % Green
+Hidden++Persona $379$ 44 $383$ $99.0$
+Hidden++No persona $393$ 44 $397$ $99.0$
+Visible++Persona $353$ 0 $353$ $100.0$
+Visible++No persona $142$ $93$ $235$ $60.4$
+$\chi^{2}(3,N=1368)=430.38$, p<.001p<.001***, $V=0.32$
+Figure 6.
+CoT Reasoning Shift: Identity-Driven vs Payoff-Optimal Keywords for Qwen Models.
+This figure compares identity-driven reasoning (short-term + social-moral keywords) against payoff-optimal reasoning (game-theoretic + strategic + long-term keywords) for Qwen 32B and Qwen 7B across all four conditions.
+Table 14.
+Pairwise Fisher exact tests for equilibrium selection among Nash outcomes in the combined Qwen models, comparing Green versus Tragedy distributions across conditions.
+Holm correction is applied across the six pairwise comparisons.
+Contrast $p$ $p_{\mathrm{Holm}}$ Effect Size (OR)
+Hidden++Persona vs. Hidden++No persona $1.000$ $1.000$ $0.96$
+Hidden++Persona vs. Visible++Persona .125.125 .376.376 $0.12$
+Hidden++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $55.33$
+Hidden++No persona vs. Visible++Persona .127.127 .376.376 $0.12$
+Hidden++No persona vs. Visible++No persona <.001<.001*** <.001<.001*** $57.38$
+Visible++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $463.89$
+Tables 13 and 14 show that the large shift in equilibrium selection for the Qwen models is concentrated in the Visible++No persona condition.
+The three remaining conditions are almost entirely Green among Nash outcomes, whereas Visible++No persona produces a much larger Tragedy share.
+This provides formal statistical support for the main-text claim that persona removal and explicit payoffs jointly shift not only whether equilibrium is reached, but also which equilibrium is selected.
+Appendix I Chain-of-Thought Analysis
+We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes.
+For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms.
+We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style.
+From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making).
+This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies.
+While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here.
+A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented below.
+Figure 7.
+Game-Theoretic Keyword Frequency Across Models and Conditions.
+This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions
+I.1. Keyword Categories
 The following keyword categories are used to analyze chain-of-thought reasoning patterns:
@@
   • Explicit information: References to “explicit”, “given”, “provided”, “shown”, “displayed”, “matrix”, “table”
-C.1. Example Chain-of-Thought Rationales
+I.2. Example Chain-of-Thought Rationales
 The following examples illustrate typical chain-of-thought reasoning patterns observed across different experimental conditions:
@@
 Example 2 (No Persona Condition, Visible Payoffs): Model: Llama-3.1-8B, Persona: Industrialist, Expected: CLEAN (Green Transition), Actual: CLEAN
+Figure 8.
+This figure compares strategic and long-term keyword frequencies across all four models in the no-persona + visible-payoffs condition (the condition enabling Tragedy equilibrium selection for Qwen models).
 “Based on the payoff matrix, I see that CLEAN has a higher maximum payoff (12.0) compared to POLLUTE (1.0).
@@
 These examples demonstrate how reasoning shifts from identity-driven considerations (Example 1 and 3) to strategic payoff optimization (Example 2) depending on persona presence and payoff visibility.
-Appendix D Supplementary Results
-Figure 2 and Tables 10, 11 and 12 contain supporting tables that provide additional detail for the findings presented in the main text.
-Table 10.
-Economic Scenario Green Transition Rates: Comparison of With Personas vs Without Personas.
-This table shows Green Transition selection rates in economic scenarios (where Tragedy is payoff-optimal).
-Each row shows a different model-variant combination.
-The table demonstrates that removing personas reduces Green Transition selection by 10-37 percentage points, indicating that personas bias models toward Green actions even in economic scenarios where Tragedy is optimal.
-This supports the finding that role identity bias prevents payoff-optimal equilibrium selection.
-Model Variant With Persona Without Persona Difference
-Llama-8B Hidden 45.0% (27/60) 26.7% (16/60) -18.3 pp
-Mistral-7B Hidden 58.3% (35/60) 21.7% (13/60) -36.6 pp
-Qwen-32B Hidden 31.7% (19/60) 15.0% (9/60) -16.7 pp
-Qwen-7B Hidden 28.3% (17/60) 18.3% (11/60) -10.0 pp
-Table 11.
-Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs.
-This table compares overall Nash equilibrium rates (across all scenarios and models) between hidden and visible payoffs conditions.
-Each row shows a different model-persona combination.
-The table reveals that with personas, hidden payoffs achieve equal or higher Nash rates than visible payoffs, suggesting reasoning scaffolding is more effective when role identity bias is present.
-Without personas, visible payoffs show large decreases for Mistral and Qwen models, but for Qwen, this is due to enabling Tragedy in economic scenarios (which is correct behavior).
-Model Persona Hidden Visible Difference
-Llama-8B Yes 77.4% (205/265) 77.0% (204/265) -0.4 pp
-Llama-8B No 75.5% (200/265) 72.1% (191/265) -3.4 pp
-Mistral-7B Yes 77.4% (205/265) 69.8% (185/265) -7.6 pp
-Qwen-32B Yes 69.1% (183/265) 57.7% (153/265) -11.4 pp
-Qwen-7B Yes 75.5% (200/265) 75.5% (200/265) 0.0 pp
-Mistral-7B No 75.8% (201/265) 38.9% (103/265) -36.9 pp
-Qwen-32B No 74.3% (197/265) 33.2% (88/265) -41.1 pp
-Qwen-7B No 75.5% (200/265) 55.5% (147/265) -20.0 pp
-Table 12.
-Persona Effect in Environmental Scenarios: Nash Equilibrium Rate Change (Without Personas vs With Personas).
-This table shows persona effects specifically for environmental scenarios.
-The structure is similar to Table 6(a), but focuses only on the 41 environmental scenarios.
-The table shows that removing personas generally reduces environmental Nash rates (negative values), with the largest effects for Mistral and Qwen models with visible payoffs (-47.8 $p$ and -58.0 $p$, respectively).
-This indicates that personas are particularly important for coordinating Green Transition in environmental scenarios, especially when explicit payoffs are provided.
-Model Hidden Visible Average
-Llama-8B -2.4 pp -6.3 pp -4.4 pp
-Mistral-7B -2.0 pp -47.8 pp -24.9 pp
-Qwen-32B +6.8 pp -58.0 pp -25.6 pp
-Qwen 7B 0.0 pp -44.9 pp -22.5 pp
-Appendix E Chain-of-Thought Analysis Figures
+Appendix J Chain-of-Thought Analysis Figures
 This section presents visualizations of chain-of-thought reasoning patterns that support the findings discussed in the main text.
 These figures illustrate how keyword frequencies shift across experimental conditions, revealing the mechanisms behind observed Nash equilibrium patterns.
-E.1. Keyword Patterns Across Conditions
+J.1. Keyword Patterns Across Conditions
 In figure 9, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning).
-This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed, and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
+This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
 Figure 9.
 CoT Keyword Frequency Patterns by Model and Condition.
-This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model, detailing the distribution of keywords within the generated rationales.
-E.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
+This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model.
+J.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
 In figure 6, the left panel (Qwen 32B) demonstrates the dramatic shift: with personas, identity-driven keywords dominate (particularly short-term thinking in hidden-payoff conditions), but without personas and with visible payoffs, payoff-optimal keywords surge dramatically, with strategic keywords appearing 368 times (unique to this condition).
@@
 This visualization directly demonstrates how removing role identity bias enables a transition from identity-aligned to payoff-optimal reasoning, correlating with the observed Nash equilibrium achievement patterns (65-90% Nash in economic scenarios for Qwen models when personas are removed and payoffs are visible).
-Figure 6.
-CoT Reasoning Shift: Identity-Driven vs Payoff-Optimal Keywords for Qwen Models.
-This figure compares identity-driven reasoning (short-term + social-moral keywords) against payoff-optimal reasoning (game-theoretic + strategic + long-term keywords) for Qwen 32B and Qwen 7B across all four conditions.
-E.3. Model Comparison of Game-Theoretic Reasoning
+J.3. Model Comparison of Game-Theoretic Reasoning
 All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition.
-However, this consistent high keyword usage does not translate to the ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas.
+However, this consistent high keyword usage does not translate to ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas.
 This suggests that while keyword frequency is informative, the combination of strategic keywords and payoff visibility is necessary for payoff-optimal equilibrium selection, which only Qwen models achieve.
-Figure 7.
-Game-Theoretic Keyword Frequency Across Models and Conditions.
-This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions.
-E.4. Strategic Keywords - A Unique Pattern in Qwen 32B
-Figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage.
+J.4. Strategic Keywords: Unique Pattern in Qwen 32B
+The figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage.
 This unique pattern corresponds to Qwen 32B’s achievement of 90% Nash equilibrium in economic scenarios and 99.7% game-theoretic reasoning in rationales.
 Long-term keywords also appear more frequently for Qwen 32B (283 mentions) compared to other models in this condition.
-This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model architectures (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.
-Figure 8.
-This figure compares strategic and long-term keyword frequencies across all four models in the no-persona + visible-payoffs condition (the condition enabling Tragedy equilibrium selection for Qwen models).
+This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model families (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.